### v1

# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

Abstract.
Hybrid storage systems (HSS) combine multiple storage devices with diverse characteristics to achieve high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which rearranges stored data across the devices to sustain high HSS performance. Prior works focus on improving only data placement or only data migration in HSS, which leads to sub-optimal HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique for HSS that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two light-weight autonomous RL agents, a data-placement agent and a data-migration agent, which adapt their policies for the current workload and HSS configuration, and coordinate with each other to improve overall HSS performance.
[…]
1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution for both high-performance and large-capacity storage requirements of data-intensive applications such as databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). HSS combine multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). A typical HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy improves HSS performance by determining the best-fit storage device in the HSS for incoming I/O requests. The data-migration policy rearranges data across the storage devices to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns.
We identify four key challenges in designing efficient data-place-ment and data-migration policies. First, in data-intensive environments, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently, which makes it hard to optimize the placement and migration policies. Second, the data-placement policy should have a low performance overhead because it operates in the critical path of I/O request handling. Third, the data-migration policy needs to migrate data across the storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the data-placement and data-migration policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
Limitations of prior works. A large body of prior work proposes various techniques to optimize either data-placement (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; ren2019archivist, ; heuristics_hyrbid_hystor_sc_2011, ; wu2012data, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data-migration (e.g., (shetti2019machine, ; doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, )) policies, but these prior works provide sub-optimal HSS performance. Our motivational study on performance-optimized (cost-optimized) HSS configurations demonstrates that even the state-of-the-art data-placement technique for HSS, Sibyl (singh2022sibyl, ), merely achieves 49.9% (41.4%) of the I/O performance of an Oracle policy that makes data-placement and data-migration decisions based on perfect knowledge of future access patterns (see §3 for more details). Based on our evaluation of various prior techniques, we identify two key reasons for such large performance gaps from the Oracle policy. First, no prior work proposes a holistic approach to optimize both data-placement and data-migration policies together. Second, heuristic- and supervised-learning-based policies (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; doudali2019kleio, ; shetti2019machine, ; ren2019archivist, ; lin2011hot, ; lin2017efficient, )) require hand-tuning to provide performance benefits and may not adapt to changes in workload characteristics and HSS states.
Since there are no prior works that focus on optimizing both data-placement and data-migration policies, we combine prior data-placement and data-migration techniques to create four extended data-management techniques (see §3 for more details). Our evaluation shows that these extended techniques underperform because their data-placement and data-migration policies are taken from two different prior approaches and are oblivious to each other’s decisions. We conclude that for high HSS performance, both placement and migration policies need to be adaptive and coordinate with each other.
Our goal is to develop a new, holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve the overall HSS performance. To this end, we propose Harmonia, [^1] Harmonia (wiki:harmonia, ) is the Greek goddess of harmony and balance, who brings peace to conflicts. a multi-agent online reinforcement learning (RL) based holistic approach for data placement and data migration in HSS.
Why RL? We choose RL for HSS data management because it offers three key benefits compared to heuristic- and supervised-learning-based approaches. First, RL dynamically adapts its decision-making to changes in workload access patterns and HSS conditions based on system-level feedback without human intervention. Second, RL can learn strategies that maximize long-term returns in complex tasks. Third, RL offers easy extensibility to various workloads and HSS configurations with minimum designer effort.
Need for multi-agent RL. Despite the benefits of an RL-based approach, a single RL agent may not effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration shows 18.8% (27.7%) lower performance on average than Sibyl on a performance-optimized (cost-optimized) HSS (see §3 for more details).
Key Idea. Harmonia is a multi-agent data-management approach in which two light-weight autonomous RL agents, a data-placement agent and a data-migration agent, improve their policies continuously, and coordinate with each other to improve HSS performance. First, we design Harmonia’s data-migration agent to identify migration candidates and their target devices by continuously monitoring previously-placed pages in HSS. Harmonia assigns a reward to the migration agent based on the impact of its migrations on future HSS performance. Second, we design Harmonia’s data-placement agent by extending the state-of-the-art data-placement technique, Sibyl (singh2022sibyl, ). To coordinate with the migration agent, we add new state features and improve reward structure for the data-placement agent. The two agents coordinate implicitly, i.e., they influence each other’s decisions through actions that affect the HSS environment. Unlike prior works (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) that propose multi-agent RL approaches for various system optimizations and use (1) identical agents, and (2) a joint-action space and shared reward functions for all agents, Harmonia is the first multi-agent RL technique to co-optimize the different tasks of data placement and migration in HSS using heterogeneous agents with independent action spaces and reward structures.
[…]
  • We demonstrate on a real Linux-based system with various HSS configurations that prior data-management approaches fall short of an Oracle policy due to the lack of combined optimization of data-placement and data-migration policies.
  • We propose Harmonia, a multi-agent RL-based data-management technique for HSS that optimizes data-placement and data-migration policies using two light-weight RL agents and achieves coordination between the agents to improve HSS performance.
  • We conduct rigorous evaluation of Harmonia on a real system with various HSS configurations of up to four devices, showing that Harmonia consistently outperforms state-of-the-art data-management techniques across a wide range of workloads.
2. Background
2.1. Hybrid Storage Systems
Fig. 1 presents an overview of a hybrid storage system (HSS). An HSS has two key components: (1) multiple storage devices of varying characteristics (e.g., I/O latency, I/O throughput/bandwidth, and device capacity) and storage protocols (e.g., NVM Express (NVMe) (nvme, ) or SATA (sata, )), and (2) an HSS management layer that orchestrates data placement and data migration in the HSS.
Figure 1. Overview of a hybrid storage system.
A typical HSS combines fast storage devices with lower capacities (i.e., high-end device in Fig. 1) and slower storage devices with larger capacities (i.e., mid-range and low-end devices in Fig. 1).
The HSS management layer has three key functions. First, it (1) provides a unified flat virtual address space for the applications to store data across all the HSS devices, (2) maps a virtual address to a logical block address (LBA) in one of the underlying storage devices, and (3) translates application I/O requests to the underlying storage devices using NVMe (nvme, ) or SATA (sata, ) protocols. The HSS management layer is similar to the md (mdlinux, ) feature in Linux, which provides software RAID functionality (RAID 0) when multiple storage devices are present. Second, the HSS management layer improves performance by (1) placing data (i.e., data placement) in the best-fit storage device for each incoming I/O request, and (2) rearranging data (i.e., data migration) across storage devices to handle changes in data access patterns after data placement and manage the faster devices effectively. Prior works (e.g., (singh2022sibyl, ; matsui2017design, ; ren2019archivist, )) demonstrate that data-placement and data-migration policies have a significant impact on HSS performance. Third, the HSS management layer stores the following metadata in the host DRAM: (1) virtual address to LBA mapping, (2) I/O request characteristics (e.g., request type, request size), and (3) storage device characteristics (e.g., device capacities and their current utilization). The HSS management layer is typically implemented in the kernel space of the host operating system.
2.2. Reinforcement Learning
Reinforcement learning (RL) (sutton_2018, ) is a machine learning (ML) approach in which an agent learns the optimal policy for a specific objective by interacting with its environment. Fig. 2 presents the overview of an RL system.
Figure 2. Overview of reinforcement learning.
An RL approach is formulated using four key components.
1. State is the representation of the environment. Assuming $S$ is the set of all possible states, $s_{t}$in $S$ is the state at time step $t$.
2. Action. At each time step, the agent observes the current state $s_{t}$ and performs an action $a_{t}$ from the set of all actions $A$. Based on the action, the environment transitions to a new state $s_{t+1}$.
3. Reward. In response to $a_{t}$ that changes environment state from $s_{t}$ to $s_{t+1}$, the agent receives a numerical reward $r_{t+1}$.
4. Policy. An agent uses policy $\pi$ to determine its actions in a given state and aims to find the policy that maximizes cumulative rewards. The optimal policy $\pi^{*}$ is determined by computing the optimal action-value function $Q^{*}$, also called the Q-value. The Q-value of a state-action pair, denoted as $Q(S,A)$, reflects the predicted cumulative reward obtained by taking action A in state S.
The agent makes decisions and learns to maximize the overall cumulative reward or return over time. While the agent may not maximize the benefit of each action, it aims to maximize the long-term consequences of its actions.
2.3. Why RL for Data Management in HSS?
Prior work (singh2022sibyl, ) shows that RL outperforms heuristic- and supervised-learning-based approaches for data placement in HSS. RL is a good fit for HSS data management because of three key reasons.
First, unlike supervised learning approaches (e.g.,  (doudali2019kleio, ; shetti2019machine, )), an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback. To demonstrate the dynamic changes in workload access patterns, we show the virtual addresses accessed and I/O request sizes during a 300-second execution of an example workload, ssd-02 (see §5 for evaluation methodology), in Figs. 3(a) and 3(b) respectively.
Figure 3. Virtual addresses accessed (top) and request sizes (bottom) captured during a 300-second execution of ssd-02 workload. Each virtual address accessed is represented as a green dot in the top figure.
We make two key observations: (1) the application accesses a broad range of virtual addresses ($1\times 10^{6}$ $1\times 10^{9}$) without fixed patterns. (2) the I/O request sizes vary from random accesses (i.e., 4KiB) to large sequential accesses (i.e., 128KiB). We observe such dynamic access pattern variations in all our chosen workloads. HSS system conditions can also change dynamically. For example, the device capacity utilization changes with each write operation, and access latencies vary significantly when multiple applications access a device concurrently. Prior heuristic- and supervised-learning-based approaches (see S3.1) make fixed placement decisions and perform migrations at infrequent intervals, and hence cannot adapt efficiently to the dynamic variations in workload access patterns and system conditions (see §3.2 for our motivational results).
Second, RL can handle complex tasks such as data placement and data migration where each action changes the environment (e.g., device capacity utilization, bandwidth utilization, garbage collection) and impacts future outcomes (e.g., I/O request latencies). Unlike supervised learning approaches, RL can learn strategies that maximize long-term returns (e.g., (sutton_2018, ; sutton1999policy, ; ke2018modeling, )). Third, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort.
We conclude that RL is a suitable approach to handle data management in HSS.
3. Motivation: Experimental Study on
Existing HSS Data-Management Techniques
3.1. Methodology
We evaluate eight prior data-management techniques (see Table 1) on a real Linux-based system with two HSS configurations, performance-optimized and cost-optimized, on 15 data-intensive workloads from three benchmark suites (see Table 5). We explain below the prior techniques evaluated, focusing on their key mechanisms. For more details on our evaluation methodology, see §5.
Table 1. Comparison of HSS Data-Management Techniques.
Technique
Data
Placement
Data
Migration
Key Details
CDE (matsui2017design, ) Heuristic Heuristic
Both policies are hand-tuned.
Data eviction during
I/O request handling.
RNN-HSS (doudali2019kleio, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
K-SVM (shetti2019machine, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
Sibyl (singh2022sibyl, ) RL Heuristic
Data eviction during
I/O request handling.
Sibyl (singh2022sibyl, )
+ K-SVM (shetti2019machine, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
Sibyl (singh2022sibyl, )
+ RNN-HSS (doudali2019kleio, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
CDE (matsui2017design, )
+ RL-Migr
Heuristic RL
Hand-tuned data-placement policy.
SAPM RL
Single RL agent for both tasks.
Harmonia RL RL
Two coordinating RL agents.
Data-Placement Techniques. We evaluate two state-of-the-art data-placement techniques for HSS: CDE (Cold Data Eviction) (matsui2017design, ) and Sibyl (singh2022sibyl, ). CDE is a heuristic-based data-placement technique that places hot or random (cold or sequential) data in the fast (slow) storage device, by determining the hotness and randomness of incoming data based on past access frequency and request size, respectively. Sibyl (singh2022sibyl, ) uses an RL agent to place data in the best-fit storage device in the HSS based on multiple features of I/O requests (e.g., request type and size) and HSS conditions (e.g., device capacity and I/O latency). When the fast device becomes full, both techniques evict the least-recently-used (LRU) data before placing a new I/O request data to the fast device.
Data-Migration Techniques. We evaluate two state-of-the-art data-migration techniques for HSS, K-SVM (shetti2019machine, ) and RNN-HSS (doudali2019kleio, ), which have three aspects in common. First, they both leverage supervised learning to identify the target storage device for migration of each stored page within the HSS based on the page’s access frequency. Second, they periodically determine migration candidates (i.e., pages) and perform data migration at fixed intervals (e.g., migration is performed after every 1000 incoming requests) in the critical path of I/O request handling. Third, they perform data placement for incoming I/O requests using a fixed heuristic policy (e.g., K-SVM places incoming requests in the fast device until its capacity reaches 100% (shetti2019machine, )). The key difference between the two techniques is in the supervised-learning method used; K-SVM uses a k-means-assisted support vector machine (SVM) to classify stored pages as belonging to a storage device based on their past access frequencies; RNN-HSS uses recurrent neural networks (RNN) to predict a page’s future access frequency, and migrate the page to its target storage device.
Extended Data-Management Techniques. To our knowledge, there is no existing technique that optimizes both data placement and data migration together, so we evaluate four extended techniques to explore the design space for a holistic data-management technique. For the first two techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, we simply combine Sibyl, the current best-performing data-placement technique, with the prior data-migration techniques explained above.
For the next two techniques, we explore the performance of an RL-based data-migration technique when combined with data placement. We design the third extended data-management technique, CDE+RL-Migr, [^2] The RL-based data-migration technique in CDE+RL-Migr is a contribution of this work. We design this technique as there are no prior RL-based data-migration techniques for HSS. by combining CDE, a heuristic data-placement technique, with an RL-based data-migration approach. The RL-based data-migration approach uses the same design as the data-migration agent proposed in Harmonia, including the state features and the reward structure (See §4.2).
In the fourth extended technique, SAPM (Single Agent for Placement and Migration), [^3] SAPM is a contribution of this work. We design SAPM to analyze the effectiveness of a single RL agent that performs both data placement and data migration. we design a single RL agent to perform both data placement and data migration. SAPM performs a placement action for every incoming I/O request, and focuses on data migration only during system idle times to avoid interference with the placement task (which is on the critical path of I/O request handling). We provide the same state features used in Harmonia (See §4.2) to SAPM and tune the hyperparameters to achieve the best performance possible with a single RL agent. We provide SAPM’s RL agent (1) an immediate reward for every placement action (based on the reward of Harmonia’s data-placement agent in Equation 2), and (2) a deferred reward for the migration actions (based on the reward structure of Harmonia’s data-migration agent in Equation 3). We provide a detailed analysis in §3.3 on why SAPM cannot effectively optimize two different tasks using a single RL agent.
3.2. Effectiveness of Prior Techniques
In this section, we conduct an experimental study to understand the effectiveness and limitations of existing data-placement and data-migration techniques for HSS. Fig. 4 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare the average request latencies of the prior techniques against two ideal approaches, Fast-Only and Oracle. First, in Fast-Only, we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and data-migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. All values in Fig. 4 are normalized to Fast-Only.
Figure 4. Performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM and Oracle on performance-optimized (top) and cost-optimized (bottom) HSS configurations. Performance is shown as average I/O latency normalized to Fast-Only. Lower is better.
We make five key observations from Fig. 4. First, all prior techniques exhibit significant performance gaps from Oracle for every tested HSS configuration and workload. The best-performing technique, Sibyl, merely achieves 49.9% (41.4%) of Oracle’s average performance in performance-optimized (cost-optimized) HSS, which strongly implies the need for holistic data management in HSS. Sibyl’s RL-based data-placement technique can adapt to workload and system changes, but its heuristic data-eviction policy on the critical path lowers its performance.
Second, Sibyl outperforms all other prior techniques, improving the average performance over CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%), respectively, in the performance-optimized (cost-optimized) HSS. CDE, K-SVM and RNN-HSS show lower performance than Sibyl because of three key reasons: (1) they use heuristic- or supervised-learning-based approaches, which cannot easily adapt to changes in workload access patterns and HSS conditions, (2) CDE performs data eviction on the critical path of I/O request handling, which contributes to 59.2% (95.7%) of the overall latency in performance-optimized (cost-optimized) HSS, and (3) RNN-HSS and K-SVM perform migrations at fixed intervals, which adds a significant overhead to I/O request tail latencies (See §6).
Third, two extended data-management techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, underperform compared to Sibyl. A naive combination of prior approaches performs poorly as the placement and migration policies in such a combination are oblivious to each other’s decisions. This leads to (1) frequent page movement across storage devices by the two policies, and (2) delayed data-placement policy convergence.
Fourth, CDE+RL-Migr shows lower performance than Sibyl by 4.6% (15.6%) on average in performance-optimized (cost-optimized) HSS. CDE+RL-Migr shows comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) because the RL-based migration agent proactively migrates frequently-read data to the fast device, which Sibyl does not do. However, CDE+RL-Migr shows lower performance in write-intensive workloads (i.e., SYSTOR17) because CDE’s heuristic data-placement policy places most incoming data in the fast storage device without considering device characteristics, resulting in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive data-placement policy.
Fifth, SAPM shows lower performance compared to Sibyl by 18.8% (27.7%) on average in the performance-optimized (cost-optimized) HSS), even though SAPM uses RL for both data placement and data migration. SAPM’s limited performance is due to two key reasons. (1) SAPM’s single RL agent is unable to learn an optimal policy for data placement and data migration concurrently, and (2) SAPM cannot identify migration candidates in a timely manner because it focuses on the data-migration task only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.3.
3.3. Need for Multi-Agent RL for
HSS Data Management
We discuss how data placement and data migration are two different tasks, and describe the limitations of a single-agent RL approach in optimizing different tasks.
3.3.1. Differences between Data Placement and Data Migration.
Data placement and data migration are two different tasks in HSS data management because of three key reasons. First, while data placement writes data from the current I/O request to the best-fit storage device, data migration monitors all previously placed data to identify candidate pages for migration and their target storage devices. Second, the objective of data placement is to minimize the I/O latency for each incoming request, but data migration aims to improve the long-term HSS performance. Third, data placement is in the critical path of I/O handling and it is performed synchronously with application writes or updates. Data migration is a background process and it is performed asynchronously to the application’s I/O requests. To describe the need for data migration in addition to data placement, we provide two examples: (1) If a page currently placed in the fast device is rarely accessed, migrating it to the slow device can free up space for more frequently-accessed pages and improve the overall HSS performance. (2) If a page on the slow device is accessed frequently (due to changes in access patterns), it should be migrated to the fast device to reduce its access latency.
3.3.2. Limitations of a Single RL Agent in Learning Diverse Tasks.
We describe two limitations of a single RL agent in optimizing multiple diverse tasks. First, a single RL agent cannot effectively learn optimal policies for multiple diverse tasks. A single-agent RL approach to optimize multiple tasks relies on task similarity (ammar2014automated, ; carroll2005task, ) to transfer the learning gained from one task to another. Task similarity refers to the resemblance of multiple tasks in their inputs, objectives, actions and rewards. Data-placement and data-migration tasks have different inputs and objectives (See §3.3.1). Prior works (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ; ammar2014automated, ; carroll2005task, ) show that a single RL agent cannot effectively learn optimal policies for multiple tasks with low task similarity because transferring the learning from one task to a dissimilar task can inhibit performance. Second, a single RL agent cannot perform two tasks concurrently. Since it performs data placement in the critical path of I/O request handling, a single agent cannot concurrently identify migration candidates, which impacts the timeliness of data migration. In §3.2, we quantitatively demonstrate that a single RL agent (SAPM) shows limited performance when handling both data placement and data migration.
We conclude that a single RL agent cannot effectively optimize both data placement and data migration in HSS, which motivates the need for a multi-agent RL approach.
3.4. Our Goal
Based on our observations and analyses in §3.2 and §3.3, we draw four conclusions. First, a holistic data-management technique is required to maximally utilize the potential of HSS. Second, RL-based techniques significantly improve performance over heuristic- and supervised-learning-based approaches. Third, given the low effectiveness of the naive combinations of existing techniques, the holistic data-management technique needs to coordinate data placement and data migration efficiently. Fourth, despite the high potential of RL for HSS data management, a single-agent-based approach has limited capability to optimize data placement and data migration simultaneously.
Our goal is to develop a new, holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve the overall HSS performance.
4. Harmonia
We design Harmonia as an online multi-agent RL technique for data management in HSS, which employs two autonomous RL agents: a data-placement agent and a data-migration agent. We implement Harmonia in the HSS management layer in the host operating system. In this section, we first describe the challenges of designing a multi-agent RL-based technique for HSS data management, and then explain Harmonia’s design and implementation in detail.
4.1. Challenges of Designing a Multi-Agent Technique for HSS
Designing a multi-agent RL-based data-management technique for HSS poses three key challenges. First, we need to design two autonomous RL agents to optimize the two different tasks of data placement and data migration (see §3.3.1). Second, we need to carefully formulate the reward structures of the two RL agents to enable coordination between the two agents (see §4.3) and improve HSS performance. Third, a multi-agent RL technique can increase computational complexity and storage overhead compared to a single-agent RL technique (see §4.6). We design Harmonia to address these challenges effectively.
While multi-agent RL itself is not a fundamentally new approach to system optimization, these prior techniques (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) cannot be adapted directly for data management in HSS because they use (1) identical RL agents that focus on a single objective, and (2) joint-action space and shared reward functions for all RL agents.
4.2. Harmonia: RL Formulation
Fig. 5 shows the overview of Harmonia’s multi-agent RL framework. The key goal of the data-placement agent is to place incoming I/O requests in the best-fit storage device in the HSS, and that of the data-migration agent is to identify migration candidates from previously-placed pages in the HSS and determine their target storage device. We design the state, action, and reward of the two agents to enable coordination between them and maximize the overall HSS performance.
Figure 5. RL Agents in Harmonia.
State. Table 2 summarizes the state features for Harmonia’s two RL agents. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics (e.g., access frequency, migration interval) when making their decisions. Request type ($req\_type$) indicates a read or a write operation. Request size ($req\_size$) is the number of 4KiB pages in the current I/O request. Access interval ($acc\_intr$) is the number of page accesses between two accesses to the same page, while access frequency ($acc\_freq$) is the total number of accesses to the page. Fast-device capacity ($fast\_cap$) is the remaining free space in the fast storage device. Current device ($curr\_dev$) indicates the storage device where the page currently resides. Migration interval ($migr\_intr$) is the number of migrations performed between two migrations of the current page. The input features are represented as a 7-dimensional tuple (observation vector) for both agents:
(1) $O_{t}=(req\_size_{t},req\_type_{t},acc\_intr_{t},acc\_freq_{t},\\ fast\_cap_{t},curr\_dev_{t},migr\_intr_{t}).$
The size of the observation vector directly impacts Harmonia’s storage and computational overheads (see §4.6), so we optimize the observation vector size in two ways. First, we carefully select a limited set of highly relevant features. Second, we group the possible values of each input feature into a small number of bins (Table 2). For features with continuous values, we choose the number of bins based on characterization of evaluated workloads. For example, the request size can fall into any of eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB, and so on) requiring three bits to represent this feature. In total, we use 32 bits for encoding the state representation.
Table 2. State features used by Harmonia
Feature Description # of bins Encoding (bits)
$req\_type$ Request type (read/write) 2 1
$req\_size$ Request size (in pages) 8 3
$acc\_intr$
Access interval
of the requested page
64 8
$acc\_freq$
Access frequency
of the requested page
64 8
$fast\_cap$
Free space
in the fast storage device
8 3
$curr\_dev$
Storage device where the
requested page currently resides
2 1
$migr\_intr$
Migration interval of a page
64 8
Action. The action space of Harmonia’s data-placement and data-migration agents depends on the number of storage devices in the HSS. The data-placement agent receives the state features for each incoming write request, and its action is to determine the best-fit storage device to place the I/O request data. The data-migration agent continuously looks at previously-placed pages in the HSS, and uses their state features to determine their target device for migration. Therefore, the two agents do not perform their actions in synchronization. In an HSS with two devices, (1) the data-placement agent can place data in either the fast or the slow device, and (2) the data-migration agent can migrate data to the fast or the slow device. The action sets of both agents are (1) small, which results in less computational complexity, and (2) easily extensible to multiple storage devices. We use a 4-bit encoding for the action to allow extensibility to more storage devices.
Reward. The reward structure is key to finding an optimal policy in RL. We design the reward structures of Harmonia’s data-placement and data-migration agents based on two key factors. First, since we focus on improving HSS performance, we use I/O request latencies in our reward structure for both agents. I/O request latencies capture the internal state (e.g., queueing delays, buffer dependencies, bandwidth utilization, effects of garbage collection, read/write latencies, non-blocking operations, error handling latencies) of storage devices in the HSS. Second, we design the reward structures to enable coordination between the two agents.
For every data-placement action at time step $t$, the data-placement agent receives a reward $R_{placement}$ from the environment at time step $t+1$. The reward for the placement agent is inversely proportional to the I/O request latency:
(2) $R_{placement}=\frac{1}{L_{t}}$
where $L_{t}$ is I/O request latency at time step $t$. The reward is higher (lower) if the data is placed in the fast (slow) device. We design a simple reward for the data-placement agent because its objective is to minimize the latency of the ongoing I/O request.
We design a different reward structure for the migration agent since its objective is different from that of data placement (see §3.3.1). The data-migration agent’s objective is to maximize the long-term performance of the HSS. Unlike the immediate impact of data placement on HSS performance, the impact of data migrations may be delayed. This is due to two key reasons: (1) the application may not access the migrated data to the fast storage device immediately, and (2) the application may not immediately leverage the free space created in the fast storage device by cold data migration. Hence, we provide a delayed reward to the migration agent to capture the long-term impact of its decisions. The reward for the migration agent is:
[…]
where $t$, $n$, $L_{i}$, $P_{migr_{t}}$ and $x$ represent the current time step, the number of incoming I/O requests whose data-placement latencies are considered, the latency of I/O request $i$, a penalty based on migration and access intervals, and the number of pages migrated before providing a reward, respectively. This reward is based on the average data-placement latencies (i.e., $\frac{\sum L_{i}}{n}$) for $n$ I/O requests that arrive after $x$ migration candidates are moved to their target devices, and serves as feedback for the migration decisions made previously. To avoid repeated migrations of the same page (ping-pong migrations), we add a small penalty $P_{migr_{t}}$ whose value is inversely proportional to the average migration and access intervals (See Table 2) of $x$ pages. Based on empirical analysis (see §6.1 for more details), we set the value of $n$ to 50 incoming requests and $x$ to 10, the size of the migration queue. We use a half-precision floating-point (16-bit) representation for the reward structure of both agents to reduce the storage overhead of training data. Our design of the reward structures of the two agents addresses the second key challenge described in §4.1.
4.3. Coordination Between Harmonia’s Agents
We achieve implicit coordination between the data-placement and data-migration agents through (1) their rewards, and (2) the effects of their actions on the HSS environment. The data-placement agent’s reward (See Equation 2) encourages effective utilization of the fast device, which is possible only if the data-migration agent proactively migrates cold data to the slow device and hot data to the fast device. The data-migration agent receives a delayed reward (See Equation 3) that is based on access latencies of $n$ incoming I/O requests after moving several migration candidates to their target devices. In this way, each agent’s decisions influence the policy learned by the other agent (See second key challenge in §4.1).
Mitigating Conflicting Decisions. To prevent conflicting placement and migration decisions, which can lead to ping-pong movement of pages across storage devices, we use two key techniques. First, the data-migration agent prioritizes the monitoring of pages that have high access and migration intervals, and ignores recently placed or migrated pages (e.g., if the page’s access interval is 1, i.e., it was recently placed). Second, we add a small penalty based on access and migration intervals to the migration agent’s reward (See Equation 3) to discourage frequent page movements. These mitigation techniques prevent the agents from making conflicting decisions.
4.4. Harmonia: Design
Fig. 6 shows the key components of Harmonia. We describe these key components below.
Figure 6. Overview of Harmonia.
Data-Placement Agent  1. An RL agent that places I/O request data in the best-fit storage device in the HSS 2 by considering I/O request characteristics and current HSS conditions. The data-placement agent consists of two separate threads. First, the data-placement training thread  3 is a background process in which the agent is trained. Second, the data-placement inference thread  4 is the only foreground thread in Harmonia. In this main decision thread, the data-placement agent decides the target storage device for incoming I/O request data in the critical path.
Data-Migration Agent 5. We implement the data migration using two background threads to avoid interference with the I/O request handling in the critical path. First, the data-migration training thread  6 is used for training the agent. Second, in the data-migration inference thread  7, the agent performs three key tasks: (i) continuously monitors previously-placed pages in the HSS. (ii) determines migration candidates and their target devices. (iii) pushes the metadata of the migration candidates into the migration queue  8. A page is not a migration candidate if the target device determined by the migration agent is the same as its current device.
Migration Queue 8. A queue that stores the metadata of the migration candidates identified by the data-migration agent. Each migration queue entry stores the logical page address and the target device identifier (e.g., 0 for slow device) of each page. We configure the migration queue length to migrate the pages in a timely manner during system idle times and avoid stale migration candidates. Based on empirical analysis (see §6.1), we set the migration queue size to 10 in our experiments.
Migration Thread 9. A background process that continuously monitors the migration queue and migrates pages across devices in the HSS during system idle times. For high-intensity workloads where the system idle times can be very short, the migration thread checks if pages related to incoming read requests are already present in the migration queue. For these pages, it issues low-priority writes (nvme, ; sata, ) to immediately schedule them for migration.
Harmonia’s five-threaded design prevents the following background processes from interfering with data placement of incoming I/O requests: (1) agent training, (2) migration candidate identification, and (3) data migration across storage devices in the HSS.
4.5. Design of Harmonia’s RL Agents
Fig. 7 shows the detailed design of each RL agent. Both agents in Harmonia have a similar design using two threads (training and inference) for ease of implementation and low storage overhead.
Figure 7. Design of RL Agent in Harmonia.
Inference Thread. Harmonia’s data-placement and data-migration agents use the inference thread to make decisions while collecting experiences to use as training data. In the inference thread, the RL agent (1) observes the input features (state) a from the I/O request data and HSS, (2) uses its inference network b to determine the action c for the I/O request, (3) selects the action d that maximizes its long-term rewards, and (4) collects the state, action and reward e as experiences f in an experience buffer g.
Training Thread. In this background thread, Harmonia uses collected experiences h from the experience buffer to train the training network i and optimize the data-placement or data-migration policy. The training thread periodically (e.g., every 1000 requests) transfers the weights of the training network to the inference network for the inference network to adapt to the current workload and HSS conditions.
Experience Buffer. Experiences are tuples of $\langle State,Action,\\ Reward,NextState\rangle$ collected by the inference thread of each agent and stored in an experience buffer to use as training data. We use separate experience buffers for the two agents because they have different goals and experiences. Based on empirical analysis, we set the experience buffer size to store the most recent 1000 experiences. A larger experience buffer can capture more access patterns but comes with significant storage overhead. The experience buffers reside in the host DRAM. Harmonia uses experience replay (dqn, ; singh2022sibyl, ), where a batch of experiences is randomly sampled from the experience buffer and used as training data for the RL agent.
Training and Inference networks. We separate the RL agent’s training and inference networks to allow parallel execution of training and inference, and remove training latency overhead from I/O request handling. Training network weights are periodically copied to the inference network (after every 1000 requests) for the inference network to adapt to current workload and HSS conditions. For the training and inference networks of both agents, we use an identical simple feed-forward neural network structure (singh2022sibyl, ) with one hidden layer of 10 neurons (de1993backpropagation, ) that uses use the swish activation function (ramachandran2017searching, ). Harmonia’s data-placement and data-migration policies select the action with the maximum predicted Q-value.
RL Algorithm. We use C51, a categorical Deep Q-Network (C51, ) for both data-placement and data-migration agents to update the Q-values $Q(S,A)$ (See §2). C51 aims to learn the distribution of Q-values for each state-action pair, which helps Harmonia capture more information from the environment (harrold2022data, ; singh2022sibyl, ).
Hyper-Parameter Tuning. We improve Harmonia’s accuracy by tuning its hyper-parameters (paine2020hyperparameter, ) using cross-validation (arlot2010survey, ) of different values. Harmonia’s agents need only one-time offline hyper-parameter tuning. Table 3 shows the hyper-parameter values chosen based on empirical analysis.
Table 3. Hyper-parameters
Hyper Parameter
Design Space
Placement
Agent
Migration
Agent
Discount Factor
($\gamma$)
0-1 0.9 0.1
Learning Rate
($\alpha$)
$1e^{-5}-1e^{0}$ $1e^{-3}$ $1e^{-2}$
Exploration Rate
($\epsilon$)
0-1 0.001 0.001
Batch Size
64-256 128 256
Experience Buffer Size
10-10000 1000 1000
The discount factor ($\gamma$) determines the balance between immediate and future rewards, while the learning rate ($\alpha$) is the rate at which neural network weights are updated. The exploration rate ($\epsilon$) balances exploration and exploitation for Harmonia’s policies. Batch size determines the number of samples processed in each training iteration. Experience buffer size is the number of most recent experiences stored for training the networks.
Exploration vs. Exploitation. For every evaluated workload, Harmonia’s data-placement and data-migration agents start with no prior knowledge and make random initial decisions (explore), and use their experiences (exploit) to gradually make optimal placement and migration decisions. To balance exploration and exploitation, we use the $\epsilon$-greedy policy (tokic2011value, ): the predicted best action is selected with (1-$\epsilon$) probability, and a random action with $\epsilon$ probability.
Convergence of RL Agents in Harmonia. Figs 8(a) and  8(b) show the convergence of Harmonia’s data-placement and data-migration policies. We show the convergence in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 5) on a performance-optimized HSS.
Figure 8. Convergence of Harmonia’s (a) Data-Placement and (b) Data-Migration RL agents. Convergence is shown in terms of reduction in training loss for each RL agent. We show the first 10000 I/O requests in this figure.
We make two key observations. First, the data-placement policy converges in less than 6000 I/O requests for this workload without any prior training. Second, the data-migration policy takes longer to converge (around 9000 I/O requests) due to the complexity of the state space, and indirect and delayed reward to the data-migration agent. We conclude that the convergence of Harmonia’s agents indicates that they learn to coordinate their actions and make optimal decisions.
4.6. Overhead Analysis
In this section, we provide an analysis of Harmonia’s overhead in terms of (1) inference and training latencies, (2) storage overhead.
4.6.1. Latency Overhead.
Harmonia has low inference and training latencies because of its (1) light-weight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). Harmonia’s two RL agents each have identical training and inference networks. We perform training and inference on host CPU because the networks are light-weight, and the neural network weights fit in the on-chip caches of the CPU in our evaluated system.
[…]
4.6.2. Storage Overhead.
The storage overhead in Harmonia constitutes (1) the neural networks in the RL agents, (2) the experience buffer, (3) the migration queue, (4) a delayed reward for the migration agent, and (5) address mapping information. First, the weights of the training and inference networks are represented using a half-precision floating-point format. With 90 16-bit weights, each neural network requires $\sim$1.5 KiB of memory. We use the same network architecture for the four networks (training and inference networks for two agents), which takes a negligible 6 KiB of memory. Due to the neural networks’ negligible size, we store their weights in the CPU caches. Second, we use an experience buffer to train each RL agent. Each experience buffer is located in the DRAM, and consumes 100 KiB to store 1000 experiences. Hence, the total storage overhead of two experience buffers is 200 KiB. Overall, Harmonia’s RL agents require 206 KiB of memory. Third, the migration queue holds the logical block address (32 bits) and target device ID (4 bits) for each queue entry. We set the migration queue size to 10 in our experiments. In total, we need 50 bytes of DRAM for the migration queue. Fourth, we store the latencies of 50 incoming I/O requests temporarily to assign a delayed reward to the migration agent, which consumes  100B. Fifth, we maintain the address mapping information for the underlying storage devices (tsukada2021metadata, ) in the HSS. Harmonia requires 32 bits to store the state features for each page (see Table 2). This overhead is less than $\sim$0.8of the total storage capacity when using a 4-KiB data placement granularity.
5. Evaluation Methodology
We evaluate Harmonia on a real system with an HSS with two (dual-HSS), three (tri-HSS), and four (quad-HSS) devices. The dual-HSS configurations include: (1) performance-optimized HSS, which combines a high-end SSD (inteloptane, ) with a mid-range SSD (intels4510, ), and (2) cost-optimized HSS, which pairs the high-end SSD (inteloptane, ) with a low-end HDD (seagate, ). We implement a custom block device driver in Linux for Harmonia to interface with the underlying storage devices in the HSS. We implement Harmonia using the TF-Agents (TFAgents, ) library. We plan to open-source Harmonia’s implementation to aid future research on HSS. Table 4 provides the details of our real system and the storage devices in the HSS configurations.
Table 4. Host System and HSS Configurations
Host System
AMD Ryzen 7 2700G(amdryzen, ),
8-cores@3.5 GHz, 8$\times$64/32 KiB L1-I/D,
4 MiB L2, 8 MiB L3,
16 GiB RDIMM DDR4 2666 MHz
Storage Devices Characteristics
H: Intel Optane SSD P4800X (inteloptane, )
375 GB, PCIe 3.0 NVMe, SLC,
R/W: 2.4/2 GB/s,
random R/W: 550000/500000 IOPS
M: Intel SSD D3-S4510 (intels4510, )
1.92 TB, SATA TLC (3D),
R/W: 560/510 MB/s,
random R/W: 895000/21000 IOPS
L: Seagate HDD ST1000DM010 (seagate, )
1 TB, SATA 6Gb/s 7200 RPM
Max. Sustained Transfer Rate: 210 MB/s
LSSD: ADATA SU630 SSD  (adatasu630, )
960 GB, SATA, TLC,
R/W: 520/450 MB/s
PMEM (Emulated): Intel Optane
Persistent Memory 200 Series  (intelpmem200, )
128 GB, Memory Mode
R/W: 7.45/2.25 GB/s (256B)
HSS Configurations Devices
Performance-Optimized high-end (H) & middle-end (M)
Cost-Optimized high-end (H) & low-end (L)
HSS with PMEM Emulated PMEM (PMEM) & high-end SSD (H)
Tri-HSS
high-end (H) & middle-end (M)
& low-end (L)
Quad-HSS
high-end (H) & middle-end (M)
& low-end SSD (LSSD)
& low-end HDD (L)
Baselines. We compare Harmonia against four state-of-the-art heuristic- and machine-learning-based data-placement and migration techniques, (1) CDE (matsui2017design, ), (2) RNN-HSS (adapted from  (doudali2019kleio, )), (3) K-SVM (shetti2019machine, ), and (4) Sibyl (singh2022sibyl, ). We evaluate four extended approaches, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM to explore the design space of holistic data-management techniques. We compare Harmonia and prior approaches with two ideal approaches: (1) Fast-Only, and (2) Oracle (meswani2015heterogeneous, ) (See § 3 for details). The Oracle serves as a reference to evaluate the accuracy of Harmonia’s decisions as it makes optimal decisions based on knowledge of future access patterns.
Workloads. We select seventeen data-intensive storage workloads from SYSTOR ’17 (lee2017understanding, ), RocksDB traces (yadgar2021ssd, ), Yahoo! Cloud Serving Benchmark (YCSB) suite (cooper2010benchmarking, ), and MLPerf Storage (mlperfstoragebenchmarks, ) that are collected from real enterprise and datacenter environments. The average workload size is approximately 50000x the fast device capacity. These workloads are chosen to represent diverse I/O access patterns with different read and write ratios, I/O request sizes, and inter-request arrival times. Table 5 reports the characteristics of the chosen workloads. In our evaluation, each workload runs in a separate thread.
Table 5. Characteristics of the evaluated I/O traces
Benchmark
Suite
Traces Read %
Avg. Request
Size (KB)
Avg. Inter-request
Arrival Time ($\mu$s)
SYSTOR17  (lee2017understanding, ) LUN0 0.2 31.7 1163.9
LUN1 0.3 34.2 1864.1
LUN2 7.6 31.1 1418.9
LUN3 3.5 42.7 734.5
LUN4 0.5 26.3 823.1
RocksDB  (yadgar2021ssd, ) ssd-00 79.9 108.9 66.4
ssd-01 73.5 75.1 40.7
ssd-02 79.9 7.5 3.3
ssd-03 79.9 9.5 3.5
ssd-04 79.9 7.8 3.6
YCSB  (cooper2010benchmarking, ) YCSB-B 51.3 45.9 9.3
YCSB-C 47.6 54.6 6.5
YCSB-D 55.9 36.1 8.5
YCSB-E 52.1 46.6 9.6
YCSB-F 49.5 53.1 6.6
MLPerf Storage  (mlperfstoragebenchmarks, ) ResNet50 80.0 172.6 500.1
CosmoFlow 83.4 180.1 1023.8
To evaluate Harmonia under real-world scenarios, we generate six multi-programmed workloads by running multiple workloads concurrently (see Table 6 for constituent workloads and their characteristics). We choose these multi-programmed workloads based on three key factors: (1) combination of read-intensive, write-intensive and mixed workloads, (2) number of concurrent workloads required to increase the intensity (i.e., shorter inter-request arrival times) of I/O requests, and (3) combination of different request sizes (See Table 6 for workload characteristics). Each constituent workload is executed using a single thread. For example, mix6 runs on eight concurrent threads for the eight constituent workloads. Multi-programmed workloads typically exhibit higher I/O request intensity and varied access patterns.
Table 6. Characteristics of multi-programmed workloads
Mix
Constituent
Workloads (cooper2010benchmarking, ; lee2017understanding, ; yadgar2021ssd, )
Description
mix1 ssd-02 and LUN4
ssd-02 is read-intensive
and LUN4 is write-intensive
mix2
LUN1 and ssd-04
LUN0 is write-intensive
and ssd-04 is read-intensive
mix3
YCSB-C and YCSB-F
Both have near-equal read-write ratio
mix4
ssd-00, ssd-04,
YCSB-A and LUN0
Two read-intensive
and two write-intensive workloads
mix5
ssd-00, LUN0,
YCSB-C and YCSB-F
Read-intensive, write-intensive and
two workloads with a
near-equal read-write ratio
mix6
YCSB-B, YCSB-D, LUN0, LUN1,
LUN4, ssd-00, ssd-02, ssd03
Two with near-equal read-write ratio,
three write-intensive
and three read-intensive
6. Evaluation
6.1. Performance Analysis
Average I/O Request Latency. Figs. 9(a) and  9(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance-optimized and cost-optimized HSS configurations. Performance is shown in terms of average request latency normalized to Fast-Only.
Figure 9. Performance of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS configurations. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make four key observations. First, Harmonia consistently outperforms all the prior and extended techniques in both HSS configurations. In performance-optimized (cost-optimized) HSS, Harmonia provides an average performance improvement of 46.4% (46.7%), 47.9% (57.7%), 43.3% (51.8%), 49.5% (31.7%), 47.6% (63.7%), 49.5% (60.0%), 36.7% (51.4%) and 45.0% (53.8%) over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM respectively.
Second, Harmonia bridges the performance gap between the best-performing prior technique, Sibyl, and Oracle by 64.2% (64.3%) on average in performance-optimized (cost-optimized) HSS.
Third, Harmonia shows significant performance benefits over prior approaches in read-intensive workloads (e.g., RocksDB, MLPerf) compared to write-intensive workloads. For example, Harmonia outperforms Sibyl by 57.5% (43.2%) in read-intensive workloads and 41.0% (39.9%) in write-intensive workloads. In read-intensive workloads, unlike Sibyl, Harmonia proactively migrates frequently-read pages to the fast device to improve read performance even when there are no writes from the application. In write-intensive workloads, while both Harmonia and Sibyl use RL-based data-placement policies, Harmonia’s data-migration agent, in coordination with the data-placement agent, further improves performance by frequently migrating cold data from the fast device to accommodate new incoming requests.
Fourth, Harmonia shows higher performance gains over prior approaches (other than Sibyl) in cost-optimized HSS than in performance-optimized HSS. This is because (1) data migration in cost-optimized HSS is more expensive than in performance-optimized HSS due to the large latency gap between the two devices, (2) prior heuristic and supervised learning approaches perform data migration during I/O request handling, which adds significant latency overheads, and (3) Harmonia avoids data eviction by freeing up space in the fast device with proactive migrations and performing migrations during system idle times. Sibyl’s RL agent places only performance-critical data in the fast device, reducing the number of data evictions in the critical path.
End-to-End Throughput. Figs. 10(a) and 10(b) show the end-to-end throughput of Harmonia and other evaluated approaches on performance-optimized and cost-optimized HSS, respectively. Throughput is shown in IOPS normalized to that of Fast-Only.
Figure 10. End-to-end throughput of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS. Throughput values (in IOPS) are normalized to Fast-Only policy. Higher is better.
We make three key observations. First, Harmonia improves average end-to-end throughput by 87.8% (210.2%), 101.3% (298.2%), 98.4% (213.1%), 49.4% (156.2%), 91.7% (362.3%), 99.8% (335.5%), 61.7% (160.6%) and 87.8% (312.6%) over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM in performance-optimized (cost-optimized) HSS. Second, Harmonia achieves an average of 71.7% (62.2%) of the throughput of Oracle on performance-optimized (cost-optimized) HSS. Third, similar to our observation from Figs.  9(a) and 9(b), Harmonia’s throughput gain over Sibyl is higher in read-intensive workloads (i.e., RocksDB) than in write-intensive. This is because, unlike Sibyl, Harmonia proactively migrates data even when there are few updates from the application, resulting in lower latencies and higher throughput.
Tail Latency. Figs. 11(a) and 11(b) show the 99th and 99.99th percentile of I/O latencies (tail latency) of Harmonia and baselines on performance-optimized and cost-optimized HSS, respectively.
Figure 11. Comparison of tail latencies of Harmonia and baselines in the 99th percentile (top) and 99.99th percentile (bottom) of I/O requests from three representative workloads, LUN0 (write-intensive), ssd-00 (read-intensive) and YCSB-B (mixed), on performance-optimized HSS.
For this analysis, we choose three representative workloads, LUN0 (write-intensive), ssd-00 (read-intensive) and YCSB-B (mixed) (see §5 for workload characteristics).
We make three key observations. First, Harmonia significantly reduces tail latency compared to all prior approaches across the three workloads. Harmonia reduces the 99th (99.99th) percentile I/O latency by 32.3% (25.3%), 48.4% (45.3%), 49.3% (46.2%), 61.8% (58.5%), 77.7% (75.6%), 92.2% (89.3%), 95.1% (92.7%) and 96.7% (97.5%) on average over Sibyl, CDE, CDE+RL-Migr, SAPM, Sibyl+K-SVM, Sibyl +RNN-HSS, K-SVM and RNN-HSS on performance-optimized (cost-optimized) HSS. Harmonia has lower tail latency because, unlike prior techniques, Harmonia performs data migration during system idle times without interfering with ongoing I/O requests. Second, only ideal policies (i.e., Oracle and Fast-Only) policies show lower tail latency than Harmonia. Third, K-SVM, RNN-HSS, Sibyl+K-SVM and Sibyl+RNN-HSS show very high tail latencies because they perform a large volume of migration at fixed intervals, which adds a significant latency overhead to the ongoing I/O requests.
We conclude that Harmonia improves average I/O latency, end-to-end throughput, and tail latency across all workloads due to its holistic and coordinated data-management approach.
Performance Sensitivity to Migration Queue Size. Figure 12(a) shows the impact of migration queue size on HSS performance. Performance is shown in terms of average I/O latency normalized to Fast-Only.
We make three key observations. First, the optimal migration queue size is 10 for our evaluated workloads. This may vary slightly across different workloads and HSS configurations. Second, low queue depths (i.e., less than 5) may result in many migration candidates being dropped, which leads to performance degradation. If the migration queue size is 0, Harmonia shows poor performance because it does not perform any data migration. Third, larger-sized queues can hold more migration candidates, but the oldest candidates may become stale (i.e., past migration decisions may not correlate with current access patterns and HSS conditions), which adversely impacts HSS performance.
Performance Sensitivity to Migration Agent’s Reward. Figure 12(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3). Performance is shown in terms of average I/O request latency normalized to Fast-Only policy.
Figure 12. Performance sensitivity to (a) migration queue sizes, and (b) number of future incoming requests used in migration agent’s reward structure. Performance is shown as average I/O latency normalized to Fast-Only. Lower is better.
We make three key observations on the optimal number of requests for the delayed reward. First, the optimal number of requests is 50 in our evaluated configuration, but may vary slightly across different workloads and HSS configurations, and depends on many factors including (1) inter-request times of incoming I/O requests after the data migration, (2) workload access patterns (e.g., access frequency and reuse distance of a page), and (3) current system conditions (e.g., device capacity utilization) Second, considering fewer requests (e.g., less than 10) may not provide the right reward to the agent because the effect of data migrations is not seen immediately in the placement latencies of these few requests. Third, if we consider more I/O requests (e.g., 250) for the reward, we see lower performance because of two reasons: (i) the impact of data migration is diluted if we consider a large number of incoming requests, and (ii) I/O requests that arrive too far into the future after data migration may not see a direct impact.
Performance on HSS with Persistent Memory (PMEM) Devices. To demonstrate the Harmonia’s extensibility to emerging persistent memory technologies, we evaluate the performance of Harmonia and baselines on an HSS with a persistent memory device (intelpmem200, ) and a high-end SSD (inteloptane, ) (see Table 4 for device characteristics). We emulate the persistent memory device using RAMDisk because these emerging devices are not easily available.
Fig. 13 shows the performance of Harmonia and baselines on an HSS with PMEM device and high-end SSD. Performance is shown as average request latency normalized to Fast-Only (i.e., PMEM device can accommodate all data).
Figure 13. Performance of Harmonia and baselines on an HSS with PMEM (intelpmem200, ) and high-end SSD (inteloptane, ). Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make four key observations. First, Harmonia outperforms all prior and extended techniques on the new HSS configuration. Harmonia provides an average performance improvement of 15.6%, 35.5%, 25.7%, 18.0%, 29.2%, 37.4%, 22.6% and 22.1% over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM respectively, and achieves 80% of Oracle’s performance. Second, CDE is the best-performing prior approach in this HSS configuration because it prefers to place most data in the low-latency PMEM device. However, CDE’s frequent data evictions in the critical path lower its performance compared to Harmonia. CDE performs better in read-intensive workloads than in write-intensive workloads because read-intensive workloads consist of fewer writes and require fewer evictions. Third, in read-intensive workloads (i.e., RocksDB, MLPerf), Harmonia shows a higher performance improvement over prior approaches because it proactively migrates data to the PMEM device, which leads to more reads from the PMEM device. In write-intensive workloads, Harmonia’s benefits come from placing only performance-critical data in the PMEM device and migrating data across devices based on access patterns. Fourth, prior approaches without a proactive data-migration policy (e.g., Sibyl) have limited performance in read-intensive workloads because they rely on application updates to migrate data.
We discuss Harmonia’s applicability to emerging memory technologies in §7. We conclude that Harmonia provides significant performance benefits for HSS configurations with emerging memory technologies such as PMEM.
6.2. Multi-Programmed Workloads
Figs. 14(a) and 14(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 6) on performance- and cost-optimized HSS, respectively. Performance is shown as average I/O latency normalized to Fast-Only.
Figure 14. Performance of Harmonia and baselines on multi-programmed workloads on performance-optimized (top) and cost-optimized (bottom) HSS. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms prior approaches in all multi-programmed workloads across both HSS configurations. In performance-optimized (cost-optimized) HSS, Harmonia (i) improves performance by 38.9% (45.4%), 34.7% (44.8%), 28.4% (33.1%), 32.8% (25.3%), 46.2% (53.0%), 40.3% (47.5%), 30.6% (33.7%), and 39.0% (57.7%) over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM in performance-optimized (cost-optimized) HSS, and (ii) achieves 75% (46%) of Oracle’s performance. Second, Harmonia’s performance improves as the number of concurrent workloads increases. On eight concurrent workloads (e.g., mix6), Harmonia shows slightly higher performance gains, 39.4% (41.8%), over Sibyl compared to other multi-programmed workloads in performance-optimized (cost-optimized) HSS. Harmonia’s RL-based policies learn access pattern variations faster than prior approaches.
We conclude that Harmonia outperforms prior approaches on all multi-programmed workloads due to better adaptability to frequent variations in workload access patterns.
6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 4). We do not evaluate other baselines because they are designed for fixed HSS configuration and lack easy extensibility. Figs. 15(a) and 15(b) show the performance of Sibyl and Harmonia on a Tri-HSS and Quad-HSS respectively. Performance is shown as average request latency normalized to Fast-Only. Harmonia can be easily extended to more devices by adding (1) the remaining capacity of the new storage device to the input features, and (2) a new node in the neural network output layer to represent the new device as a target device for data placement and data migration.
Figure 15. Performance of Sibyl and Harmonia on Tri-HSS (top) and Quad-HSS (bottom) configurations. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Sibyl by 37.0% (42.0%) on average in Tri-HSS (Quad-HSS). Second, in several workloads (e.g., RocksDB, MLPerf), Harmonia shows higher performance gains (up to 58.3%) over Sibyl than the average performance improvement because Harmonia places performance-critical data in the high-end and mid-range devices and uses its data-migration policy to move cold data to slow devices during idle times. In contrast, Sibyl frequently evicts data to the low-end devices in the critical path, adding significant latency overhead.
We conclude that Harmonia demonstrates extensibility to multiple HSS configurations and achieves high performance across all tested workloads and HSS configurations.
6.4. Lifetime of Devices in HSS
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to the total data originating from the workload. A higher WA value adversely impacts the lifetime of storage devices. Figs. 16(a) and 16(b) show WA of Harmonia and baselines on performance-optimized and cost-optimized HSS, respectively.
Figure 16. Write amplification of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS. Lower is better.
We make two key observations. First, in both HSS configurations, only Sibyl and RNN-HSS have a lower WA than Harmonia. Harmonia’s average WA of 1.54 (1.55) is higher than Sibyl’s 1.49 (1.45) and RNN-HSS’s 1.08 (1.06) in performance-optimized (cost-optimized) HSS because it proactively migrates data to the fast device in read-intensive workloads to improve read performance. Second, in both HSS configurations, Harmonia has lower WA (1.24) in write-intensive workloads (i.e., SYSTOR17) compared to its WA (1.8) in read-intensive workloads (i.e., RocksDB, MLPerf). In write-intensive workloads, Harmonia learns to move data during update operations and performs fewer migrations.
Write Traffic Distribution in Harmonia. Fig.  17 shows the distribution of writes caused by Harmonia’s migration policy across fast and slow devices in performance-optimized and cost-optimized HSS.
Figure 17. Write traffic due to Harmonia’s migrations on performance-optimized (left) and cost-optimized (right) HSS.
We make two observations. First, Harmonia migrates more data from fast device to slow device in both configurations because it effectively utilizes fast device for frequently-accessed data. Second, Harmonia has more writes to fast device in read-intensive workloads (i.e., RocksDB and MLPerf) than in write-intensive workloads because it proactively migrates frequently-read data to the fast device even when there are no application updates.
We conclude that Harmonia’s adaptive policies provide higher performance at slightly higher average WA compared to the best-performing prior approach.
7. Discussion
Alternate Objectives and Reward Structures for Harmonia. Reward formulation is critical for finding optimal policies in RL. Careful design of reward structure is necessary to achieve coordination in a multi-agent system with autonomous agents. Harmonia’s focus is to improve HSS performance. Hence, we design the RL agents’ reward structure based on I/O latencies.
However, we can design Harmonia for other objectives such as improving device lifetimes, fairness, quality of service (QoS) and energy efficiency. For example, to improve storage device lifetimes, we can design the state features and reward structure to include (1) number of writes performed on each device, (2) device endurance in terms of program/erase cycles, (3) write amplification factor (WAF), and (4) device reliability properties (e.g., NAND flash memory type (SLC/TLC/QLC), ECC correction capability). To promote or disincentivize write traffic to each device, the reward should be appropriate to the device where the data is placed or migrated to. Harmonia can accommodate fairness and QoS objectives by including factors such as bandwidth utilization, device queue depths and application requirements. Combining different objectives is an interesting research direction, and we leave the exploration of these objectives and reward structures to future work.
Extensibility of Harmonia. We describe Harmonia’s extensibility to a large number of storage devices in the HSS. Harmonia, unlike prior heuristic and supervised learning based approaches, is easily scalable with minimal designer effort. When a new device is added to the HSS, we only need to extend the state features and actions of Harmonia’s RL agents. We add (1) capacity utilization of the new device as a state feature (see Table 2), and (2) an output node in each RL agent’s neural network to represent the action of placing or migrating a page to the new device. However, the larger state and action spaces resulting from scaling Harmonia to more devices leads to two challenges: (1) the data-placement and data-migration policies may take longer to converge to an optimal policy, and (2) the metadata overhead to store the state features increases.
Applicability to Other Systems. Data placement and data migration are critical for the high performance of other heterogeneous systems such as tiered storage systems, hybrid memory systems, disaggregated storage, and networked storage systems. Harmonia can be adapted to these systems, but they come with diverse constraints and requirements.
In a tiered storage system, data placement is always performed on the top-tier device, and hence Harmonia’s data-placement agent can be replaced by a heuristic approach. However, Harmonia’s data-migration agent gains significance due to data migration across tiers. We can add multiple data-migration agents to perform migration between each pair of adjacent tiers.
Hybrid memory systems have strict latency constraints, and require Harmonia to be redesigned for low-latency decision-making. Disaggregated storage systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) for large-scale memory pooling can leverage Harmonia’s benefits when deployed on CXL fabric controller. Networked storage systems with local and remote storage devices can also benefit from Harmonia’s RL framework. We hope Harmonia can assist in designing adaptive data-management policies for other heterogeneous systems.
8. Related Work
To our knowledge, Harmonia is the first work to use multi-agent RL for combined optimization of data-placement and data-migration policies in HSS. In this section, we review other related work in HSS data management, and RL in storage/memory systems.
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. Hot Random Off-loading (HRO) (lin2011hot, ) uses a 0-1 knapsack model to allocate or migrate files between the storage devices. ADLAM (zhang2010adaptive, ) uses limited workload characteristics to tune data-migration policies in cloud storage systems. Vengerov (vengerov2008reinforcement, ) uses a fuzzy rule base architecture to perform data migration in a hierarchical storage system. These works do not perform (1) real system evaluation or (2) combined optimization of placement and migration policies using multiple RL agents.
Reinforcement Learning in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) use RL to improve different aspects of storage systems. Prior works  (liu2019learning, ; wang2020reinforcement, ) use RL to improve cloud storage utilization. RLAlloc (li2023rlalloc, ) improves SSD throughput and Quality of Service (QoS). Other works (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ) use RL to optimize SSD maintenance activities (e.g., garbage collection) and error mitigation. Harmonia is orthogonal to all these works and can be used in conjunction with these techniques on an HSS.
RL for Architectural Optimizations. Prior works use RL for data prefetching (e.g., (pythia, ; peled2015semantic, )), memory scheduling (e.g., (ipek2008self, ; multi_scheduler_HPCA_2012, )), cache replacement strategies (liu2020imitation, ) and network-on-chip arbitration (rl_NOC_AIDArc_2018, ; lin2020deepNOC, ). These works provide evidence that RL holds significant potential for intelligent high-performance computer systems.
Multi-Agent RL Systems. Several works propose using multi-agent reinforcement learning (MARL) for various system optimizations (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )). Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) use MARL to improve tail latency of Serverless Function-as-a-Service (FaaS), targeting different systems and objectives than Harmonia. Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) use MARL for dynamic partitioning of last level cache and multi-level caches. Shen et al. (shen2022multiagent, ) propose MARL-assisted cache cleaning for shingled managed recording drives. These techniques have four key differences from Harmonia: (1) their objectives are different from HSS data placement and migration, (2) they use identical RL agents unlike Harmonia’s heterogeneous agents, (3) they use joint actions and shared rewards for all agents while Harmonia’s agents have independent actions, and (4) they use table-based RL unlike Harmonia’s neural network based RL. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.
9. Conclusion
We propose Harmonia, the first multi-agent reinforcement learning based holistic data-management technique for HSS that optimizes data placement and data migration together. We show that Harmonia’s design enables coordination between the data-placement and data-migration agents to improve HSS performance. Our real-system evaluation demonstrates that Harmonia significantly outperforms prior HSS data-management techniques on four HSS configurations across a wide range of workloads. Harmonia is easily extensible to different HSS configurations. The performance benefits of Harmonia come with low latency and storage overheads.