### v2

# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

Abstract.
Hybrid storage systems (HSS) combine multiple storage devices with diverse characteristics to achieve high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which rearranges stored data (i.e., prefetches hot data and evicts cold data) across the devices to sustain high HSS performance. Prior works focus on improving only data placement or only data migration in HSS, which leads to relatively low HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS, and thus significantly improve system performance. We demonstrate the need for multiple reinforcement learning (RL) agents to accomplish our goal. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two lightweight autonomous RL agents, a data-placement agent and a data-migration agent, which adapt their policies for the current workload and HSS configuration, and coordinate with each other to improve overall HSS performance.
[…]
copyright: none
1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution for high-performance and large-capacity storage requirements of data-intensive applications such as machine learning (DLRMFacebook, ; near_human_3, ; ho2020denoising, ), large language models (meta_llama, ; jiang2024mixtral, ; liu2024deepseek, ), databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). An HSS combines multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). A typical HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy determines the best-fit storage device in the HSS for incoming I/O requests. The data-migration policy rearranges data across the storage devices (i.e., prefetches frequently-accessed data to the fast device and evicts cold data to the slow device) to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns.
We identify four key challenges in designing efficient data-place-ment and data-migration policies. First, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently in data-intensive environments, which makes it hard to optimize the policies. Second, the data-placement policy should have a low performance overhead as it operates on the critical path of I/O request handling. Third, the data-migration policy needs to migrate data across storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the two policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
Limitations of prior works. Prior works propose techniques for either data placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data migration (e.g., (doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )), but they provide relatively low performance when employed together or alone. Our motivational study on performance- (cost-) optimized HSS configurations demonstrates that even the state-of-the-art HSS data-placement technique, Sibyl (singh2022sibyl, ), only achieves 49.9% (41.4%) of the performance of an Oracle policy (see §3 for details). We identify two key reasons for such large performance gaps. First, no prior work proposes a holistic approach to optimize both data placement and data migration together. Second, heuristic- and supervised-learning-based policies (e.g., (singh2022sibyl, ; matsui2017design, ; matsui2017tri, ; doudali2019kleio, ; shetti2019machine, ; ren2019archivist, ; lin2011hot, ; lin2017efficient, )) require hand-tuning and may not adapt to changes in workload characteristics and HSS states.
Since there are no prior works that optimize both placement and migration, we combine prior techniques to create four extended data-management techniques (see §3). Our evaluation shows that these extended techniques underperform because they are naive combinations of prior data-placement and data-migration approaches that are oblivious to each other’s decisions. We conclude that for high HSS performance, both data placement and migration policies need to coordinate with each other.
Our goal is to design a holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve overall system performance. To this end, we propose Harmonia, [^1] Harmonia (wiki:harmonia, ): Greek goddess of harmony and balance, who brings peace to conflicts. a multi-agent online reinforcement learning (RL) based holistic approach to co-optimize HSS data placement and data migration.
Need for multi-agent RL. We describe the benefits of RL for HSS data management in §3.1. Despite these benefits, a single RL agent cannot effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration leads to 18.8% (27.7%) lower average performance than Sibyl, which performs only data placement, on a performance- (cost-) optimized HSS (see §3.4). This is because learning to optimize data migration interferes with the process of learning an optimal data placement policy as they are two different tasks (See §3.2).
Key Idea. Harmonia employs two coordinating lightweight autonomous RL agents, respectively for data placement and migration, to improve HSS performance. First, we design Harmonia’s data-placement agent by extending the state-of-the-art placement technique, Sibyl (singh2022sibyl, ). To coordinate with the migration agent, we add new state features and improve reward structure for the placement agent. Second, we design Harmonia’s data-migration agent to identify migration candidates and their target devices by continuously monitoring previously-placed pages in HSS. Harmonia assigns rewards to the migration agent based on the impact of its migrations on future HSS performance. The two agents influence each other’s decisions through actions that affect the HSS environment.
[…]
This work makes the following key contributions:
  • We show that prior data-management approaches have low HSS performance because they do not optimize both data placement and data migration, and we quantify this on a real Linux system with various HSS configurations.
  • We propose Harmonia, a multi-agent RL-based data-management technique for HSS that co-optimizes data-placement and data-migration policies using two lightweight coordinating RL agents to improve HSS performance.
  • Via a rigorous real-system evaluation of Harmonia on various HSS configurations, we demonstrate that Harmonia consistently outperforms state-of-the-art data-management techniques across a wide range of data-intensive workloads.
2. Background
2.1. Hybrid Storage Systems
Fig. 1 presents an overview of a hybrid storage system (HSS). An HSS has two key components: (1) multiple storage devices of varying characteristics (e.g., I/O latency, I/O bandwidth, and device capacity) and storage protocols (e.g., NVM Express (NVMe) (nvme, ) or SATA (sata, )), and (2) an HSS management layer that orchestrates data placement and data migration in the HSS.
Figure 1. Overview of a hybrid storage system.
A typical HSS combines fast storage devices with lower capacities (i.e., high-end device in Fig. 1) and slower storage devices with larger capacities (i.e., mid-range and low-end devices in Fig. 1).
HSS Management Layer. The HSS management layer has three key functions. First, it (1) provides a unified flat virtual address space for the applications to store data across all HSS devices, (2) maps a virtual address to a logical block address (LBA) in one of the underlying storage devices, and (3) translates application I/O requests to the devices using NVMe (nvme, ) or SATA (sata, ) protocols. The HSS management layer is similar to the md (mdlinux, ) feature in Linux, which provides software RAID functionality (RAID 0) when multiple storage devices are present. Second, it (1) places data (i.e., data placement) in the best-fit storage device for each incoming I/O request, and (2) rearranges data (i.e., data migration) across storage devices. Prior works (e.g., (singh2022sibyl, ; matsui2017design, ; ren2019archivist, )) demonstrate that data-placement and migration policies have a significant impact on HSS performance. Third, the HSS management layer stores the following metadata in the host DRAM: (1) virtual address to LBA mapping, (2) I/O request characteristics (e.g., request type, request size), and (3) storage device characteristics (e.g., device capacities and their current utilization). The HSS management layer is typically implemented in the kernel space of the host OS.
2.2. Reinforcement Learning
Reinforcement learning (RL) (sutton_2018, ) is a machine learning (ML) approach in which an agent learns the optimal policy for a specific objective by interacting with its environment. Fig. 2 presents the overview of an RL system.
Figure 2. Overview of reinforcement learning.
An RL approach is formulated using four key components.
1. State is the representation of the environment. Assuming $S$ is the set of all possible states, $s_{t}$in $S$ is the state at time step $t$.
2. Action. At each time step, the agent observes the current state $s_{t}$ and performs an action $a_{t}$ from the set of all actions $A$. Based on the action, the environment transitions to a new state $s_{t+1}$.
3. Reward. In response to $a_{t}$ that changes environment state from $s_{t}$ to $s_{t+1}$, the agent receives a numerical reward $r_{t+1}$.
4. Policy. An agent uses policy $\pi$ to determine its actions in a given state and aims to find the policy that maximizes cumulative rewards. The optimal policy $\pi^{*}$ is determined by computing the optimal action-value function $Q^{*}$, also called the Q-value. The Q-value of a state-action pair, denoted as $Q(S,A)$, reflects the predicted cumulative reward obtained by taking action A in state S.
The agent makes decisions and learns to maximize the overall cumulative reward or return over time. While the agent may not maximize the benefit of each action, it aims to maximize the long-term consequences of its actions.
2.3. Multi-Agent Reinforcement Learning
Multi-agent RL (e.g., (canese2021multi, ; rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) is an extension of traditional RL where multiple RL agents interact with a shared environment and and with one another to learn optimal policies. Unlike RL with a single agent, multi-agent RL can effectively optimize multiple different tasks. A single RL agent cannot effectively learn optimal policies for different tasks because it relies on task similarity (ammar2014automated, ; carroll2005task, ) to transfer the learning gained from one task to another. Task similarity refers to the resemblance of tasks in their inputs, objectives, and actions. Prior works (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ; ammar2014automated, ; carroll2005task, ) show that a single RL agent cannot effectively learn optimal policies for multiple tasks with low task similarity because transferring the learning from one task to a dissimilar task can inhibit performance.
3. Motivation
In this section, we describe (1) why RL is a good fit for HSS data management, (2) the need for a multi-agent RL approach, and (3) prior approaches to HSS data management and their effectiveness.
3.1. Why RL for Data Management in HSS?
RL is a good fit for HSS data management for three key reasons.
First, unlike heuristic and supervised-learning approaches, an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback without human intervention. Prior work Sibyl (singh2022sibyl, ) shows that RL outperforms heuristic-(e.g., (matsui2017design, ; matsui2017tri, )) and supervised-learning-based (e.g.,  (doudali2019kleio, ; shetti2019machine, )) approaches for HSS data placement. However, Sibyl partially uses RL for HSS data-management (i.e., RL for data-placement and a heuristic approach for data migration) that leaves a large gap between its performance and that of an Oracle approach (as we discuss in §3.4).
Second, RL can handle complex tasks such as placement and migration where each action changes the environment (e.g., device capacity utilization, bandwidth utilization, garbage collection) and impacts future outcomes (e.g., I/O request latencies). Unlike supervised-learning approaches, RL learns strategies that maximize long-term returns (e.g., (sutton_2018, ; sutton1999policy, ; ke2018modeling, )), i.e., overall HSS performance.
Third, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort (as partially demonstrated in Sibyl (singh2022sibyl, )).
3.2. Need for Multi-Agent RL for HSS
Data placement and data migration are two different tasks in HSS, and a multi-agent RL approach is a great fit for optimizing these different tasks (see §2.3).
Data placement and data migration in HSS are two different tasks for three key reasons. First, they operate on different data in the HSS. While data placement writes data from the current I/O request to the best-fit storage device, data migration monitors previously placed data to identify candidate pages for migration and their target storage devices. Second, their objectives are different. The objective of data placement is to minimize the I/O latency for each incoming request, but data migration aims to improve long-term HSS performance. Third, these two tasks are performed asynchronously with each other. Data placement is on the critical path of I/O handling and is performed synchronously with application writes or updates. Data migration is a background process and it is performed asynchronously to the application’s I/O requests.
To describe the need for data migration in addition to data placement, we provide two examples: (1) If a page in the fast device is rarely accessed, migrating it to the slow device can free up space for more frequently-accessed pages and improve overall HSS performance. (2) If a page on the slow device is accessed frequently (due to access pattern changes), it should be migrated to the fast device to reduce its access latency. In §3.4, we quantitatively demonstrate that a single RL agent (SAPM) shows limited performance when handling both data placement and migration in HSS. We conclude that multi-agent is a great fit for optimizing both data placement and migration in HSS, which are complex and different tasks.
3.3. Prior Approaches to HSS Data Management
Our focus in this work is to co-optimize placement and migration in HSS. In this section, we briefly introduce prior HSS data-management techniques. However, there is no existing technique that optimizes both placement and migration in HSS. We discuss different approaches to optimize both placement and migration by combining (or extending) prior HSS data-management techniques.
Data-Placement Techniques. We evaluate two state-of-the-art HSS data-placement techniques: CDE (Cold Data Eviction) (matsui2017design, ) and Sibyl (singh2022sibyl, ). CDE is a heuristic data-placement technique that places hot or random (cold or sequential) data in the fast (slow) device, by determining the hotness and randomness of incoming data based on past access frequency and request size, respectively. Sibyl (singh2022sibyl, ) uses an RL agent to place data in the best-fit HSS storage device based on multiple features of I/O requests (e.g., request type and size) and HSS conditions (e.g., device capacity, I/O latency). When the fast device becomes full, both techniques evict the least-recently-used (LRU) data before placing new I/O request data in the fast device.
Data-Migration Techniques. We evaluate two state-of-the-art HSS data-migration techniques, K-SVM (shetti2019machine, ) and RNN-HSS (doudali2019kleio, ), which have three aspects in common. First, they both leverage supervised learning to identify the target device for migration of each stored page within the HSS based on the page’s access frequency. Second, they periodically determine migration candidates (i.e., pages) and perform data migration at fixed intervals (e.g., migration is performed after every 1000 incoming requests) on the critical path of I/O request handling. Third, they perform data placement for incoming I/O requests using a fixed heuristic policy (e.g., K-SVM places incoming requests in the fast device until its capacity reaches 100% (shetti2019machine, )). The key difference between the two techniques is in the supervised-learning method used; K-SVM uses a k-means-assisted support vector machine (SVM) to classify stored pages as belonging to a storage device based on their past access frequencies; RNN-HSS uses recurrent neural networks (RNN) to predict a page’s future access frequency, and migrate the page to its target device.
Extended Data-Management Techniques. Since there is no existing technique that optimizes both data placement and data migration together, we evaluate four extended techniques to explore the design space for a holistic data-management technique. For the first two techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, we simply combine Sibyl, the current best-performing data-placement technique, with the prior data-migration techniques explained above.
We design the third extended data-management technique, CDE+RL-Migr, [^2] The RL-based data-migration technique in CDE+RL-Migr is a contribution of this work. We design this as there are no prior RL-based HSS data-migration techniques. by combining CDE, a heuristic data-placement technique, with an RL-based data-migration approach. The RL-based data-migration approach uses the same design as the data-migration agent proposed in Harmonia, including the state features and the reward structure (See §4.2).
In the fourth extended technique, SAPM (Single Agent for Placement and Migration), we design a single RL agent to perform both data placement and data migration to analyze the effectiveness of a single RL agent at HSS data management. SAPM performs a placement action for every incoming I/O request, and focuses on data migration only during system idle times to avoid interference with the placement task (which is on the critical path of I/O request handling). We provide the same state features used in Harmonia (See §4.2) to SAPM and tune the hyperparameters to achieve the best performance possible with a single RL agent. We provide SAPM’s RL agent (1) an immediate reward for every placement action (based on the reward of Harmonia’s data-placement agent in Equation 2), and (2) a deferred reward for the migration actions (based on the reward structure of Harmonia’s data-migration agent in Equation 3). We provide background in §2.3 on why a single RL agent cannot effectively optimize two different tasks.
3.4. Effectiveness of Prior Techniques
Methodology. We evaluate eight prior data-management techniques on a real system with two HSS configurations, performance- and cost-optimized, on 17 data-intensive workloads from 3 benchmark suites (see Table 4). See §5 for our evaluation methodology.
Performance Results. Fig. 3 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare their average request latencies against two ideal approaches, Fast-Only and Oracle. First, in Fast-Only, we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. All values in Fig. 3 are normalized to Fast-Only.
Figure 3. Performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM and Oracle on performance-optimized (top) and cost-optimized (bottom) HSS configurations. Performance is shown as average I/O latency normalized to Fast-Only. Lower is better.
We make five key observations from Fig. 3. First, all prior techniques exhibit significant performance gaps from Oracle for every tested HSS and workload. The best-performing technique, Sibyl, only achieves 49.9% (41.4%) of Oracle’s performance in performance- (cost-) optimized HSS, which strongly implies the need for holistic data management in HSS. Sibyl’s RL-based data-placement technique can adapt to workload and system changes, but its heuristic data-eviction policy on the critical path lowers its performance.
Second, Sibyl outperforms all other prior techniques, improving the average performance over CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%), respectively, in the performance- (cost-) optimized HSS. Sibyl outperforms CDE, K-SVM and RNN-HSS because they use heuristic- or supervised-learning-based approaches, which cannot easily adapt to changes in workload access patterns and HSS conditions.
Third, Sibyl+K-SVM and Sibyl+RNN-HSS underperform compared to Sibyl. A naive combination of prior approaches performs poorly as the placement and migration policies are oblivious to each other’s decisions. This leads to (1) conflicting decisions from the two policies, and (2) delayed data-placement policy convergence.
Fourth, CDE+RL-Migr shows lower performance than Sibyl by 4.6% (15.6%) on average in performance- (cost-) optimized HSS. CDE+RL-Migr shows comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) because the RL-based migration agent proactively prefetches frequently-read data to the fast device, which Sibyl does not do. However, it underperforms in write-intensive workloads (i.e., SYSTOR17) because CDE’s heuristic data-placement policy places most incoming data in the fast storage device without considering device characteristics, resulting in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive placement policy.
Fifth, SAPM shows lower performance compared to Sibyl by 18.8% (27.7%) on average in the performance- (cost-) optimized HSS, even though SAPM uses RL for both data placement and migration. SAPM’s limited performance is due to two key reasons. (1) SAPM’s single RL agent is unable to learn an optimal policy for data placement and data migration concurrently, and (2) SAPM cannot identify migration candidates in a timely manner because it focuses on the data-migration task only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.2.
3.5. Our Goal
Based on our observations and analyses in §3.1, §3.2, §3.3, and §3.4, we conclude that a multi-agent RL approach’s ability to optimize different tasks concurrently alleviates the limitations of a single-agent RL approach.
Our goal is to design a holistic HSS data-management technique that optimizes both data-placement and data-migration policies using a multi-agent RL approach to improve the overall HSS performance.
4. Harmonia
We design Harmonia as an online multi-agent RL technique for HSS data management with two autonomous RL agents: a data-placement agent and a data-migration agent. We implement Harmonia in the HSS management layer in the host operating system. In this section, we first describe the challenges of designing a multi-agent RL-based technique for HSS data management, and then explain Harmonia’s design and implementation in detail.
4.1. Challenges of Designing a Multi-Agent HSS Technique
Designing a multi-agent RL-based HSS data-management technique poses three key challenges. First, we need to design two autonomous RL agents to optimize the different tasks of data placement and migration (see §3.2). Second, we need to formulate the reward structures of the two RL agents to enable coordination between them (see §4.3) and improve HSS performance. Third, having multiple RL agents can increase computational complexity and storage overhead (see §4.6). While multi-agent RL itself is not a fundamentally new approach to system optimization, prior techniques (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) cannot be adapted directly for HSS data management because they use (1) identical RL agents with a single objective, and (2) joint-action space and shared reward functions for all RL agents.
4.2. Harmonia: RL Formulation
Fig. 4 shows the overview of Harmonia’s multi-agent RL framework. The key goal of the data-placement agent is to place incoming I/O requests in the best-fit storage device, and that of the data-migration agent is to identify migration candidates from previously-placed pages and determine their target storage device. We design the state, action, and reward of the two agents to enable coordination between them and maximize the overall HSS performance.
Figure 4. RL Agents in Harmonia.
State. Table 1 summarizes the state features for Harmonia’s two RL agents. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics when making their decisions. The state features are represented as a 7-dimensional tuple (observation vector) for both agents:
(1) $O_{t}=(req\_size_{t},req\_type_{t},acc\_intr_{t},acc\_freq_{t},\\ fast\_cap_{t},curr\_dev_{t},migr\_intr_{t}).$
The observation vector size directly impacts Harmonia’s storage and computational overheads (see §4.6), so we (1) carefully select a limited set of highly relevant features, and (2) group state feature values into a small number of bins (Table 1) based on workload characterization. For example, request size can fall into eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB) requiring three bits to represent. In total, we use 32 bits for encoding the state representation.
Table 1. State features used by Harmonia
Feature Description # of bins Encoding (bits)
$req\_type$ Request type (read/write) 2 1
$req\_size$ Request size (in pages) 8 3
$acc\_intr$
Access interval
of the requested page
64 8
$acc\_freq$
Access frequency
of the requested page
64 8
$fast\_cap$
Free space
in the fast storage device
8 3
$curr\_dev$
Storage device where the
requested page currently resides
2 1
$migr\_intr$
Migration interval of a page
64 8
Action. The action space of Harmonia’s agents depends on the number of devices in the HSS. The data-placement agent receives the state features of incoming write requests, and its action is to determine the best-fit device to place request data. The data-migration agent continuously looks at the state features of previously-placed pages to determine their target device for migration. Therefore, the two agents do not perform their actions in synchronization. In an HSS with two devices, (1) the data-placement agent can place data in the fast or the slow device, and (2) the data-migration agent can migrate data to the fast or the slow device. The action sets of both agents are small and extensible, resulting in less computational complexity and easier multi-device scalability. We encode actions with 4 bits, allowing extensibility to more storage devices.
Reward. The reward structure is key to finding an optimal policy in RL. We design the reward structures of Harmonia’s agents based on two key factors. First, since we focus on HSS performance, we use I/O request latencies in our reward structure for both agents. I/O request latencies capture the internal state (e.g., queueing delays, buffer dependencies, bandwidth utilization, effects of garbage collection, read/write latencies, non-blocking operations, error handling latencies) of storage devices in the HSS. Second, we design the reward structures to enable coordination between the two agents.
For every placement action at time step $t$, the data-placement agent receives a reward $R_{placement}$ from the environment at time step $t+1$ that is inversely proportional to the I/O request latency:
(2) $R_{placement}=\frac{1}{L_{t}}$
where $L_{t}$ is I/O request latency at time step $t$. The reward is higher (lower) if the data is placed in the fast (slow) device. We design a simple reward for the data-placement agent because its objective is to minimize the latency of the ongoing I/O request.
The data-migration agent’s objective is to maximize long-term HSS performance, which is different from the placement agent’s objective (see §3.2). Unlike the immediate impact of data placement on HSS performance, the impact of data migrations (i.e., prefetching frequently-accessed data to the fast device and evicting cold data to the slow device) may be delayed due to two key reasons: (1) the application may not access the prefetched data immediately, and (2) the application may not immediately leverage the free space created in the fast device by cold data migration. Hence, we provide a delayed reward to the migration agent to capture the long-term impact of its decisions. The reward for the migration agent is:
[…]
where $t$ is the current time step, $n$ is the number of incoming I/O requests whose data-placement latencies are considered, $L_{i}$ is the latency of I/O request $i$, $P_{migr_{t}}$ is a penalty based on migration and access intervals, and $x$ is the number of pages migrated before providing a reward. This reward is based on the average data-placement latencies (i.e., $\frac{\sum L_{i}}{n}$) for $n$ I/O requests that arrive after $x$ migration candidates are moved to their target devices, and serves as feedback for the migration decisions made previously. To avoid repeated migrations of the same page (ping-pong migrations), we add a small penalty $P_{migr_{t}}$ whose value is inversely proportional to the average migration and access intervals (See Table 1) of $x$ pages. Based on empirical analysis (see §6.1), we set the value of $n$ to 50 and $x$ to 10, the size of the migration queue. We use a half-precision floating-point (16-bit) representation for the reward structure of both agents to reduce the storage overhead of training data.
4.3. Coordination Between Harmonia’s Agents
We achieve coordination between the two agents through (1) their rewards, and (2) the effects of their actions on the HSS environment. The placement agent’s reward (See Equation 2) encourages effective utilization of the fast device, which is possible only if the migration agent proactively migrates cold (hot) data to the slow (fast) device. The data-migration agent receives a delayed reward (See Equation 3) that is based on access latencies of $n$ incoming I/O requests after migrating several candidate pages. In this way, each agent’s decisions influence the policy learned by the other agent.
Mitigating Conflicting Decisions. To prevent conflicting placement and migration decisions, we use two key techniques. First, the migration agent prioritizes the monitoring of pages with high access and migration intervals, and ignores recently placed or migrated pages (e.g., if the page’s access interval is 1, i.e., it was recently placed). Second, we add a small penalty based on access and migration intervals to the migration agent’s reward (See Equation 3) to discourage frequent page movements.
4.4. Harmonia: Design
Fig. 5 shows the key components of Harmonia. We describe these key components below.
Figure 5. Overview of Harmonia.
Data-Placement Agent  1. An RL agent that places I/O request data in the best-fit storage device in the HSS 2 based on I/O request characteristics and HSS conditions. This agent consists of two threads. First, the data-placement training thread  3 is a background process in which the agent is trained. Second, the data-placement inference thread  4 is Harmonia’s only foreground thread. In this main decision thread, the data-placement agent decides the target storage device for incoming I/O request data on the critical path.
Data-Migration Agent 5. We implement the data-migration agent using two background threads to avoid interference with the I/O request handling on the critical path. First, the data-migration training thread  6 is used for training the agent. Second, in the data-migration inference thread  7, the agent performs three key tasks: (i) continuously monitors previously-placed pages in the HSS. (ii) determines migration candidates and their target devices. (iii) pushes the metadata of the migration candidates into the migration queue  8. A page is not a migration candidate if the target device determined by the migration agent is the same as its current device.
Migration Queue 8. A queue that stores the logical page address and the target device identifier (e.g., 0 for slow device) of each migration candidate identified by the migration agent. Based on empirical analysis (see §6.1), we set the migration queue size to 10 in our experiments to migrate pages in a timely manner during system idle times and avoid stale migration candidates.
Migration Thread 9. A background process that continuously monitors the migration queue and migrates pages across devices during system idle times. Harmonia samples device bandwidth at frequent intervals using real-time tools (e.g., iostat) to perform migrations. For high-intensity workloads with short system idle times, the migration thread checks if pages related to incoming read requests are already in the migration queue and issues low-priority writes (nvme, ; sata, ) to immediately schedule them for migration.
Harmonia’s five-threaded design prevents the following background processes from interfering with data placement of incoming I/O requests: (1) agent training, (2) migration candidate identification, and (3) data migration across storage devices in the HSS.
4.5. Design of Harmonia’s RL Agents
Fig. 6 shows the detailed design of each RL agent. Both agents in Harmonia have a similar design using two threads (training and inference) for ease of implementation and low storage overhead.
Figure 6. Design of RL Agent in Harmonia.
Inference Thread. Harmonia’s RL agents use the inference thread to make decisions while collecting experiences to use as training data. In this thread, the agent (1) observes the input features (state) a from the I/O request data and HSS, (2) performs inference using its inference network b, (3) selects the action c that maximizes d its long-term returns, and (4) collects the state, action and reward e as experiences f in an experience buffer g.
Training Thread. In this background thread, Harmonia uses collected experiences h from the experience buffer to train the training network i and optimize the placement or migration policy.
Experience Buffer. Experiences are tuples of $\langle State,Action,\\ Reward,NextState\rangle$ collected by the inference thread of each agent and stored in an experience buffer to use as training data. We use separate experience buffers for the two agents because they have different goals and experiences. Based on empirical analysis, we set the experience buffer size to store the most recent 1000 experiences. A larger experience buffer can capture more access patterns but comes with significant storage overhead. The experience buffers reside in the host DRAM. Harmonia uses experience replay (dqn, ; singh2022sibyl, ), where a batch of experiences is randomly sampled from the experience buffer and used as training data for the RL agent.
Training and Inference networks. We separate the RL agent’s training and inference networks to allow parallel execution of training and inference, and remove training latency overhead from I/O request handling. Training network weights are periodically copied to the inference network (after every 1000 requests) for the inference network to adapt to current workload and HSS conditions. For the training and inference networks of both agents, we use an identical simple feed-forward neural network structure (singh2022sibyl, ) with one hidden layer of 10 neurons (de1993backpropagation, ) that uses use the swish activation function (ramachandran2017searching, ). Harmonia’s data-placement and data-migration policies select the action with the maximum predicted Q-value.
RL Algorithm. We use C51, a categorical Deep Q-Network (C51, ) for both data-placement and data-migration agents to update the Q-values $Q(S,A)$ (See §2). C51 aims to learn the distribution of Q-values for each state-action pair, which helps Harmonia capture more information from the environment (harrold2022data, ; singh2022sibyl, ).
Hyper-Parameter Tuning. We improve Harmonia’s accuracy by tuning the hyper-parameters (paine2020hyperparameter, ) of each agent using cross-validation (arlot2010survey, ) of different values. Harmonia’s agents need only one-time offline hyper-parameter tuning. Table 2 shows the hyper-parameter values chosen based on empirical analysis.
Table 2. Hyper-parameters
Hyper Parameter
Design Space
Placement
Agent
Migration
Agent
Discount Factor
($\gamma$)
0-1 0.9 0.1
Learning Rate
($\alpha$)
$1e^{-5}-1e^{0}$ $1e^{-3}$ $1e^{-2}$
Exploration Rate
($\epsilon$)
0-1 0.001 0.001
Batch Size
64-256 128 256
Experience Buffer Size
10-10000 1000 1000
The discount factor ($\gamma$) determines the balance between immediate and future rewards, while the learning rate ($\alpha$) is the rate at which neural network weights are updated. The exploration rate ($\epsilon$) balances exploration and exploitation for Harmonia’s policies. Batch size determines the number of samples processed in each training iteration. Experience buffer size is the number of most recent experiences stored for training the networks.
Exploration vs. Exploitation. Harmonia’s RL agents start with no prior knowledge of the workload or HSS, and make random initial decisions (explore), and use their experiences (exploit) to gradually make optimal decisions. To balance exploration and exploitation, we use the $\epsilon$-greedy policy (tokic2011value, ): the predicted best action is selected with (1-$\epsilon$) probability, and a random action with $\epsilon$ probability.
Convergence of RL Agents in Harmonia. Figs 7(a) and  7(b) show the convergence of Harmonia’s data-placement and data-migration policies in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 4) on a performance-optimized HSS.
Figure 7. Convergence of Harmonia’s (a) Data-Placement and (b) Data-Migration RL agents shown in terms of reduction in training loss. We show only the first 10000 I/O requests.
We make two key observations: (1) The placement policy converges in less than 6000 I/O requests without prior training. (2) The migration policy takes longer to converge (around 9000 I/O requests) due to indirect and delayed reward to the migration agent. The convergence of Harmonia’s policies indicates that the agents learn to coordinate their actions and make optimal decisions.
4.6. Overhead Analysis
In this section, we provide an analysis of Harmonia’s overhead in terms of (1) inference and training latencies, (2) storage overhead.
4.6.1. Latency Overhead.
Harmonia has low inference and training latencies because of its (1) lightweight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). Harmonia’s two RL agents have identical networks. We perform training and inference on host CPU as the neural networks are lightweight, and their weights fit in the CPU on-chip caches in our evaluated system.
[…]
4.6.2. Storage Overhead.
Harmonia’s storage overhead includes (1) RL agents’ neural networks, (2) experience buffers, (3) migration queue, (4) migration agent’s reward, and (5) address mapping information. First, neural network weights are represented using a half-precision floating-point format. With 90 16-bit weights, each network requires $\sim$1.5 KiB of memory. The four networks (training and inference networks for two agents) take a negligible 6 KiB of memory, so we store their weights in the CPU caches. Second, to train each RL agent, we use an experience buffer located in the DRAM that consumes 100 KiB to store 1000 experiences. Hence, the total storage overhead of two experience buffers is 200 KiB. Overall, Harmonia’s RL agents require 206 KiB of memory. Third, the migration queue holds the logical block address (32 bits) and target device ID (4 bits) for each queue entry. The migration queue size is 10 in our experiments, requiring 50 bytes of DRAM. Fourth, we store the latencies of 50 incoming I/O requests temporarily to assign a delayed reward to the migration agent, which consumes  100B. Fifth, we maintain address mapping information for HSS devices (tsukada2021metadata, ). Harmonia requires 32 bits to store the state features for each page (see Table 1). This overhead is less than $\sim$0.8of the total storage capacity when using a 4-KiB data placement granularity.
5. Evaluation Methodology
We evaluate Harmonia on a real system with an HSS with two (dual-HSS), three (tri-HSS), and four (quad-HSS) devices. The dual-HSS configurations include: (1) performance-optimized HSS, which combines a high-end SSD (inteloptane, ) with a mid-range SSD (intels4510, ), and (2) cost-optimized HSS, which pairs the high-end SSD (inteloptane, ) with a low-end HDD (seagate, ). We implement a custom block device driver in Linux for Harmonia to interface with HSS devices. We implement Harmonia using the TF-Agents (TFAgents, ) library. Table 3 describes our real system and the HSS configurations.
Table 3. Host System and HSS Configurations
Host System
AMD Ryzen 7 2700G(amdryzen, ),
8-cores@3.5 GHz, 8$\times$64/32 KiB L1-I/D,
4 MiB L2, 8 MiB L3,
16 GiB RDIMM DDR4 2666 MHz
Storage Devices Characteristics
H: Intel Optane SSD P4800X (inteloptane, )
375 GB, PCIe 3.0 NVMe, SLC,
R/W: 2.4/2 GB/s,
random R/W: 550000/500000 IOPS
M: Intel SSD D3-S4510 (intels4510, )
1.92 TB, SATA TLC (3D),
R/W: 560/510 MB/s,
random R/W: 895000/21000 IOPS
L: Seagate HDD ST1000DM010 (seagate, )
1 TB, SATA 6Gb/s 7200 RPM
Max. Sustained Transfer Rate: 210 MB/s
LSSD: ADATA SU630 SSD  (adatasu630, )
960 GB, SATA, TLC,
R/W: 520/450 MB/s
PMEM (Emulated): Intel Optane
Persistent Memory 200 Series  (intelpmem200, )
128 GB, Memory Mode
R/W: 7.45/2.25 GB/s (256B)
HSS Configurations Devices
Performance-Optimized high-end (H) & middle-end (M)
Cost-Optimized high-end (H) & low-end (L)
HSS with PMEM Emulated PMEM (PMEM) & high-end SSD (H)
Tri-HSS
high-end (H) & middle-end (M) & low-end (L)
Quad-HSS
high-end (H) & middle-end (M) &
low-end SSD (LSSD) & low-end HDD (L)
Baselines. We compare Harmonia against four state-of-the-art heuristic- and machine-learning-based data-placement and migration techniques, (1) CDE (matsui2017design, ), (2) RNN-HSS (adapted from  (doudali2019kleio, )), (3) K-SVM (shetti2019machine, ), and (4) Sibyl (singh2022sibyl, ). We evaluate four extended approaches, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM to explore the design space of holistic data-management techniques. We also compare against two ideal approaches: (1) Fast-Only, and (2) Oracle (meswani2015heterogeneous, ) (See § 3). Oracle serves as a reference to evaluate the accuracy of Harmonia’s decisions as it makes optimal decisions based on knowledge of future access patterns.
Workloads. We select seventeen data-intensive storage workloads from SYSTOR ’17 (lee2017understanding, ), RocksDB traces (yadgar2021ssd, ), Yahoo! Cloud Serving Benchmark (YCSB) suite (cooper2010benchmarking, ), and MLPerf Storage [^3] We generated the traces by running MLPerf applications on our evaluated system. (mlperfstoragebenchmarks, ) from real enterprise and datacenter environments. The average workload size is approximately 50000x the fast device capacity. We choose these workloads to represent diverse I/O access patterns with different read and write ratios, I/O request sizes, and inter-request times. Table 4 reports the characteristics of the chosen workloads. In our evaluation, each workload runs in a separate thread.
Table 4. Characteristics of the evaluated I/O traces
Benchmark
Suite
Traces Read %
Avg. Request
Size (KB)
Avg. Inter-
Request Time ($\mu$s)
SYSTOR17  (lee2017understanding, ) LUN0 0.2 31.7 1163.9
LUN1 0.3 34.2 1864.1
LUN2 7.6 31.1 1418.9
LUN3 3.5 42.7 734.5
LUN4 0.5 26.3 823.1
RocksDB  (yadgar2021ssd, ) ssd-00 79.9 108.9 66.4
ssd-01 73.5 75.1 40.7
ssd-02 79.9 7.5 3.3
ssd-03 79.9 9.5 3.5
ssd-04 79.9 7.8 3.6
YCSB  (cooper2010benchmarking, ) YCSB-B 51.3 45.9 9.3
YCSB-C 47.6 54.6 6.5
YCSB-D 55.9 36.1 8.5
YCSB-E 52.1 46.6 9.6
YCSB-F 49.5 53.1 6.6
MLPerf Storage  (mlperfstoragebenchmarks, ) ResNet50 80.0 172.6 500.1
CosmoFlow 83.4 180.1 1023.8
To evaluate Harmonia under real-world scenarios, we generate six multi-programmed workloads by running multiple workloads concurrently (see Table 5). We choose these multi-programmed workloads based on three key factors: (1) combination of read-intensive, write-intensive and mixed workloads, (2) number of concurrent workloads required for high I/O intensity (i.e., shorter inter-request times), and (3) combination of different request sizes. Each constituent workload is executed using a single thread. For example, mix6 runs on eight concurrent threads for the eight constituent workloads. Multi-programmed workloads typically exhibit higher I/O request intensity and varied access patterns.
Table 5. Characteristics of multi-programmed workloads
Mix
Constituent
Workloads (cooper2010benchmarking, ; lee2017understanding, ; yadgar2021ssd, )
Description
mix1 ssd-02 and LUN4
ssd-02 is read-intensive
and LUN4 is write-intensive
mix2
LUN1 and ssd-04
LUN0 is write-intensive
and ssd-04 is read-intensive
mix3
YCSB-C and YCSB-F
Both have near-equal read-write ratio
mix4
ssd-00, ssd-04,
YCSB-A and LUN0
Two read-intensive
and two write-intensive workloads
mix5
ssd-00, LUN0,
YCSB-C and YCSB-F
Read-intensive, write-intensive and
two workloads with a
near-equal read-write ratio
mix6
YCSB-B, YCSB-D, LUN0, LUN1,
LUN4, ssd-00, ssd-02, ssd03
Two with near-equal read-write ratio,
three write-intensive
and three read-intensive
6. Evaluation
6.1. Performance Analysis
Average I/O Request Latency and End-to-End Throughput. Figs. 8(a) and  8(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance- and cost-optimized HSS configurations. Figs. 9(a) and 9(b) show their end-to-end throughput on performance- and cost-optimized HSS, respectively.
Figure 8. Performance of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS configurations, shown as average request latency normalized to Fast-Only. Lower is better.
Figure 9. End-to-end throughput of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS. Throughput values (in IOPS) are normalized to Fast-Only policy. Higher is better.
We make four key observations. First, Harmonia consistently outperforms all baselines in both HSS configurations. In performance- (cost-) optimized HSS, Harmonia improves performance by 49.5% (31.7%) and end-to-end throughput by 49.4% (156.2%) over the best-performing prior approach, Sibyl, on average.
Second, Harmonia bridges the performance gap between Sibyl and Oracle by 64.2% (64.3%) and achieves 71.7% (62.2%) of Oracle’s throughput on average in performance- (cost-) optimized HSS.
Third, Harmonia shows higher performance and throughput benefits over baselines in read-intensive workloads (e.g., RocksDB, MLPerf) compared to write-intensive workloads. For example, Harmonia outperforms Sibyl by 57.5% (43.2%) in read-intensive workloads and 41.0% (39.9%) in write-intensive workloads. In read-intensive workloads, unlike Sibyl, Harmonia proactively prefetches frequently-read pages to the fast device to improve read performance even when there are no writes from the application, resulting in lower latencies and higher throughput. In write-intensive workloads, while both Harmonia and Sibyl use RL-based placement, Harmonia’s migration agent, in coordination with the placement agent, further improves performance by frequently migrating cold data from the fast device to accommodate new incoming requests.
Fourth, Harmonia shows higher performance and throughput gains over baselines (other than Sibyl) in cost-optimized than in performance-optimized HSS. This is because (1) data migration in cost-optimized HSS is more expensive than in performance-optimized HSS due to the large latency gap between the two devices, (2) prior approaches perform migration during I/O request handling, which adds significant latency overheads, and (3) Harmonia frees up space in the fast device with proactive migrations and performing migrations during system idle times. Sibyl’s RL agent places only performance-critical data in the fast device, reducing the number of data evictions on the critical path.
Tail Latency. Figs. 10(a) and 10(b) show the 99th and 99.99th percentile I/O latencies (tail latency) of Harmonia and baselines on performance-optimized HSS.
Figure 10. Comparison of tail latencies of Harmonia and baselines in the 99th percentile (top) and 99.99th percentile (bottom) of I/O requests from three representative workloads, LUN0 (write-intensive), ssd-00 (read-intensive) and YCSB-B (mixed), on performance-optimized HSS.
For this analysis, we choose three representative workloads, LUN0 (write-intensive), ssd-00 (read-intensive) and YCSB-B (mixed) (see §5 for workload characteristics).
We make two key observations. First, Harmonia significantly reduces tail latency compared to all baselines across the three workloads. Harmonia reduces the 99th (99.99th) percentile I/O latency by 32.3% (25.3%) on average over Sibyl. Harmonia has lower tail latency because, unlike prior techniques, Harmonia performs data migration during system idle times to avoid interference with ongoing I/O requests. Second, K-SVM, RNN-HSS, Sibyl+K-SVM and Sibyl+RNN-HSS show very high tail latencies because they perform a large volume of migration at fixed intervals, which adds a significant latency overhead to the ongoing I/O requests.
We conclude that Harmonia improves average I/O latency, end-to-end throughput, and tail latency across all workloads due to its holistic and coordinated data-management approach.
Performance Sensitivity to Migration Queue Size. Fig.  11(a) shows the impact of migration queue size on HSS performance.
We make three key observations. First, optimal migration queue size is 10 for our evaluated workloads. This may vary slightly across workloads and HSS configurations. Second, low queue depths (i.e., ¡ 5) may result in many migration candidates being dropped, which leads to performance degradation. If the migration queue size is 0, Harmonia shows poor performance because it does not perform any data migration. Third, larger-sized queues can hold more migration candidates, but the oldest candidates may become stale (i.e., past migration decisions may not correlate with current access patterns and HSS conditions), which adversely impacts HSS performance.
Performance Sensitivity to Migration Agent’s Reward. Fig.  11(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3).
Figure 11. Performance sensitivity to (a) migration queue sizes, and (b) number of future incoming requests used in migration agent’s reward structure. Performance is shown as average I/O latency normalized to Fast-Only. Lower is better.
We make three key observations. First, the optimal number of requests is 50 in our evaluated configuration, but may vary slightly across workloads and HSS configurations, and depends on factors including (1) inter-request times of incoming I/O requests after migration, (2) workload access patterns (e.g., access frequency and reuse distance of a page), and (3) system conditions (e.g., device capacity utilization). Second, considering fewer requests (e.g., ¡ 10) may not provide the right reward to the agent because the effect of migrations is not seen immediately in the placement latencies of these few requests. Third, if we consider more I/O requests (e.g., 250), we see lower performance because of two reasons: (i) the impact of data migration is diluted if we consider a large number of incoming requests, and (ii) I/O requests that arrive too far into the future after data migration may not see a direct impact.
Performance on HSS with Persistent Memory (PMEM) Devices. To demonstrate the Harmonia’s extensibility to emerging persistent memory technologies, we evaluate Harmonia and baselines on an HSS with a persistent memory device (intelpmem200, ) and a high-end SSD (inteloptane, ) (see Table 3). We emulate the PMEM device using RAMDisk because these emerging devices are not easily available. Fig. 12 shows the performance of Harmonia and baselines on HSS with PMEM. Performance is shown as average request latency normalized to Fast-Only (i.e., PMEM device can store all data).
Figure 12. Performance of Harmonia and baselines on an HSS with PMEM (intelpmem200, ) and high-end SSD (inteloptane, ) shown as average request latency normalized to Fast-Only. Lower is better.
We make three key observations. First, Harmonia outperforms the best-performing baseline, CDE, by 15.6% and achieves 80% of Oracle’s performance. Second, CDE outperforms other baselines in this HSS configuration as it places most data in the low-latency PMEM device. However, CDE’s frequent data evictions on the critical path lower its performance compared to Harmonia. CDE performs better in read-intensive workloads because they consist of fewer writes and require fewer evictions. Third, in read-intensive workloads (i.e., RocksDB, MLPerf), Harmonia shows a higher performance improvement as it proactively prefetches data to the PMEM device, which leads to more reads from the PMEM device. We discuss Harmonia’s applicability to emerging memory technologies in §7.
6.2. Multi-Programmed Workloads
Figs. 13(a) and 13(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 5) on performance- and cost-optimized HSS, respectively.
Figure 13. Performance of Harmonia and baselines on multi-programmed workloads on performance-optimized (top) and cost-optimized (bottom) HSS. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms prior approaches in all multi-programmed workloads across both HSS configurations. In performance- (cost-) optimized HSS, Harmonia (i) outperforms Sibyl by 32.8% (25.3%), and (ii) achieves 75% (46%) of Oracle’s performance. Second, Harmonia’s performance improves as the number of concurrent workloads increases. On eight concurrent workloads (e.g., mix6), Harmonia shows higher performance gains, 39.4% (41.8%), over Sibyl compared to other multi-programmed workloads. Harmonia’s RL-based policies learn access pattern variations faster than prior approaches.
6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 3). We do not evaluate other baselines as they are designed for fixed HSS configurations and lack easy extensibility. Figs. 14(a) and 14(b) show the performance of Sibyl and Harmonia on a Tri- and Quad-HSS. We discuss Harmonia’s extensibility in §7.
Figure 14. Performance of Sibyl and Harmonia on Tri-HSS (top) and Quad-HSS (bottom). Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Sibyl by 37.0% (42.0%) on average in Tri-HSS (Quad-HSS) as Harmonia places performance-critical data in the high-end and mid-range devices and uses its migration policy to move cold data to slow devices during idle times. In contrast, Sibyl frequently evicts data to the low-end devices on the critical path, adding significant latency overhead. Second, in read-intensive workloads (e.g., RocksDB, MLPerf), Harmonia shows higher performance gains (up to 58.3%) over Sibyl because Sibyl lacks a migration policy and relies on application updates to migrate data across devices.
6.4. Lifetime of Devices in HSS
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to data originating from the workload. A higher WA adversely impacts device lifetimes. Figs. 15(a) and 15(b) show WA of Harmonia and baselines on performance- and cost-optimized HSS, respectively.
Figure 15. Write amplification of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS. Lower is better.
We make two key observations. First, only Sibyl and RNN-HSS have a lower WA than Harmonia. Harmonia’s average WA of 1.54 (1.55) is higher than Sibyl’s 1.49 (1.45) and RNN-HSS’s 1.08 (1.06) in performance- (cost-) optimized HSS as it proactively prefetches data to the fast device in read-intensive workloads to improve read performance. Second, in both HSS configurations, Harmonia has lower WA (1.24) in write-intensive workloads (i.e., SYSTOR17) compared to its WA (1.8) in read-intensive workloads (i.e., RocksDB, MLPerf). In write-intensive workloads, Harmonia learns to move data during update operations and performs fewer migrations.
Write Traffic Distribution in Harmonia. Fig.  16 shows the write distribution of Harmonia’s migration policy across devices in performance- and cost-optimized HSS.
Figure 16. Write traffic due to Harmonia’s migrations on performance-optimized (left) and cost-optimized (right) HSS.
We make two observations. First, Harmonia migrates more data to slow device as it effectively utilizes fast device for frequently-accessed data. Second, Harmonia has more writes to fast device in read-intensive workloads (i.e., RocksDB, MLPerf) as it proactively prefetches frequently-read data to the fast device even when there are no application updates.
We conclude that Harmonia’s adaptive policies provide higher performance at slightly higher average WA compared to Sibyl.
7. Discussion
Alternate Objectives and Reward Structures. Since Harmonia’s focus is to improve HSS performance, we design the RL agents’ reward structure based on I/O latencies. We can design Harmonia for other objectives such as device lifetimes, fairness, quality of service (QoS) and energy efficiency. For example, to improve device lifetimes, the state features and reward structure can include (1) writes performed on each device, (2) device endurance in terms of program/erase cycles, (3) write amplification (WA), and (4) device reliability properties (e.g., NAND flash type (SLC/TLC/QLC), ECC capability). To manage write traffic to each device, the reward should be appropriate to the device where the data is placed or migrated. Harmonia can accommodate fairness and QoS objectives by including factors such as bandwidth utilization, device queue depths and application requirements. We leave the exploration of these objectives and reward structures to future work.
Extensibility of Harmonia. Harmonia is easily extensible to a large number of storage devices with minimal designer effort. When a new device is added to the HSS, we only add (1) capacity utilization of the new device as a state feature (see Table 1), and (2) an output node in each RL agent’s neural network to represent the action of placing or migrating a page to the new device. However, scaling Harmonia to more devices can lead to two challenges: (1) the large state and action spaces can delay the convergence to an optimal policy, and (2) the metadata overhead increases.
Applicability to Other Systems. Harmonia can be adapted to other heterogeneous systems such as tiered storage, hybrid memory, disaggregated storage and networked storage, which have diverse constraints and requirements. In a tiered storage system, placement is always performed on the top-tier device, and hence Harmonia’s placement agent can be replaced by a heuristic approach. The migration agent gains significance due to data migration across tiers. We can add multiple migration agents for each pair of adjacent tiers. Hybrid memory systems have strict latency constraints, and require Harmonia to be redesigned for low-latency decision-making. Disaggregated storage systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) for large-scale memory pooling can leverage Harmonia’s benefits when deployed on CXL fabric controller. We hope Harmonia can assist in designing adaptive policies for other heterogeneous systems.
8. Related Work
In this section, we review other related work in HSS data management, and RL in storage/memory systems.
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. Hot Random Off-loading (HRO) (lin2011hot, ) uses a 0-1 knapsack model to allocate or migrate files between the storage devices. Vengerov (vengerov2008reinforcement, ) uses a fuzzy rule base architecture to perform data migration in a hierarchical storage system. These works do not perform (1) real system evaluation or (2) combined optimization of placement and migration policies using multiple RL agents.
RL in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) use RL to improve different aspects of storage systems. Prior works  (liu2019learning, ; wang2020reinforcement, ) use RL to improve cloud storage utilization. RLAlloc (li2023rlalloc, ) improves SSD throughput and Quality of Service (QoS). Other works (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ) use RL to optimize SSD maintenance activities (e.g., garbage collection) and error mitigation. Harmonia is orthogonal to all these works and can be used in conjunction with these techniques on an HSS.
Multi-Agent RL Systems. Several works propose using multi-agent RL (MARL) for various system optimizations (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )). Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) use MARL to improve tail latency of Serverless Function-as-a-Service (FaaS), targeting different systems and objectives than Harmonia. Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) use MARL for dynamic partitioning of last level cache and multi-level caches. Shen et al. (shen2022multiagent, ) propose MARL-assisted cache cleaning for shingled managed recording drives. These techniques have four key differences from Harmonia: (1) their objectives are different from HSS data placement and migration, (2) they use identical RL agents unlike Harmonia’s heterogeneous agents, (3) they use joint actions and shared rewards for all agents, and (4) they use table-based RL unlike Harmonia’s neural network based RL. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.
9. Conclusion
We propose Harmonia, the first multi-agent reinforcement learning based holistic data-management technique for HSS that optimizes data placement and data migration together. We show that Harmonia’s design enables coordination between the data-placement and data-migration agents to improve HSS performance. Our real-system evaluation demonstrates that Harmonia significantly outperforms prior HSS data-management techniques on four HSS configurations across a wide range of workloads. Harmonia is easily extensible to different HSS configurations. The performance benefits of Harmonia come with low latency and storage overheads.