### v3

# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

Abstract.
Hybrid storage systems (HSS) integrate multiple storage devices with diverse characteristics to deliver high performance and capacity at low cost. The performance of an HSS highly depends on the effectiveness of two key policies: (1) the data-placement policy, which determines the best-fit storage device for incoming data, and (2) the data-migration policy, which dynamically rearranges stored data (i.e., prefetches hot data and evicts cold data) across the devices to sustain high HSS performance. Prior works optimize either data placement or data migration in isolation, which leads to suboptimal HSS performance. Unfortunately, no prior work tries to optimize both policies together.
Our goal is to design a holistic data-management technique that optimizes both data-placement and data-migration policies to fully exploit the potential of an HSS, and thus significantly improve system performance. We propose Harmonia, a multi-agent reinforcement learning (RL)-based data-management technique that employs two lightweight autonomous RL agents, a data-placement agent and a data-migration agent, that adapt their policies for the current workload and HSS configuration while coordinating with each other to improve overall HSS performance.
We evaluate Harmonia on real HSS configurations with up to four heterogeneous storage devices and seventeen data-intensive workloads. On performance-optimized (cost-optimized) HSS with two storage devices, Harmonia outperforms the best-performing prior approach by 49.5% (31.7%) on average. On an HSS with three (four) devices, Harmonia outperforms the best-performing prior work by 37.0% (42.0%) on average. Harmonia’s performance benefits come with low latency (240nsns for inference) and storage overheads (206 KiB in DRAM for both RL agents combined). We will open-source Harmonia’s implementation to aid future research on HSS.
1. Introduction
A hybrid storage system (HSS) (xiao2016hs, ; wang2017larger, ; lee2014mining, ; felter2011reliability, ; bu2012optimization, ; krish2016efficient, ; lin2011hot, ; niu2018hybrid, ; tai2015sla, ; ou2014edm, ; cheng2015amc, ; zong2014faststor, ; matsui2017design, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; kim2011hybridstore, ; liu2019hierarchical, ; luo2020optimal, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ) is a cost-effective solution to meet the high-performance and large-capacity storage requirements of data-intensive applications such as machine learning (DLRMFacebook, ; near_human_3, ; ho2020denoising, ), large language models (meta_llama, ; jiang2024mixtral, ; liu2024deepseek, ), databases (chan-signmod-1998, ; FastBitA9, ; redis-bitmaps, ), graph processing (beamer-SC-2012, ), and genome analysis (lander-nature-2001, ; altschul-jmb-1990, ; myers-jacm-1999, ). An HSS combines multiple storage devices (e.g., solid-state drives (SSDs) (inteloptane, ; intels4510, ; samsung2017znand, ; samsung-980pro, ; samsungmlc, ), hard disk drives (HDDs)) of diverse characteristics (e.g., I/O bandwidth, storage capacity). Typically, an HSS consists of fast storage devices with limited capacities and slow storage devices with large capacities.
The performance of an HSS is highly dependent on two key policies, data placement and data migration. The data-placement policy determines the most suitable (i.e., best-fit) storage device in the HSS for each incoming I/O request. The data-migration policy rearranges data across the storage devices (i.e., prefetches frequently-accessed data to the fast device and evicts cold data to the slow device) to sustain high HSS performance, which can degrade over time due to (1) misplacement of data, and (2) changes in workload access patterns and HSS conditions.
We identify four key challenges in designing effective data-placement and data-migration policies. First, workload access patterns and HSS conditions (e.g., access latencies, device capacity utilization) can change frequently in data-intensive environments, which makes it hard to optimize the policies. Second, the data-placement policy operates on the critical path of I/O request handling and it should have a low performance overhead. Third, the data-migration policy needs to migrate data across storage devices in a timely manner without impacting the latency of incoming I/O requests. Fourth, the two policies should not make conflicting decisions, which can adversely impact HSS performance and device lifetimes.
Limitations of prior works. Prior works propose techniques to optimize either data placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, ; wu2019data, ; yang2014statistics, ; yolchuyev2021data, ; strunk2012hybrid, )) or data migration (e.g., (doudali2019kleio, ; lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) in isolation, which results in poor performance. For example, an optimized data-placement policy cannot sustain long-term performance if the data-migration decisions remain static or oblivious to workload changes. Similarly, an optimized migration policy cannot compensate for sub-optimal placement decisions that cause unnecessary data movement, wasted bandwidth, and increased device wear. Our motivational study on performance- (cost-) optimized HSS configurations demonstrates that even the state-of-the-art HSS data-placement technique, Sibyl (singh2022sibyl, ), only achieves 49.9% (41.4%) of the performance of an Oracle policy (see §3 for details). While Sibyl optimizes its data-placement policy to adapt to workload and system changes, the large performance gap is due to Sibyl’s heuristic data-migration policy that (ii) considers a limited number of workload features (e.g., access frequency of pages) to identify migration data, and cannot adapt to changes in workload characteristics and HSS states, and (ii) performs migration in the critical path of I/O handling, which adds a significant latency overhead.
Since no prior work jointly optimizes both placement and migration, we combine state-of-the-art data-placement and data-migration techniques to create four extended data-management techniques (see §3). Our evaluation shows that these extended techniques underperform compared to their constituent techniques because the policies operate independently, which leads to conflicting decisions, unnecessary data movement, and wasted system resources. We conclude that a naive combination of data placement and migration policies provides sub-optimal performance, and we need to co-design both the policies with coordination to achieve high HSS performance.
Our goal is to design a holistic data-management technique for HSS that optimizes both data-placement and data-migration policies in a coordinated way to improve overall system performance. To this end, we propose Harmonia, [^1] Harmonia (wiki:harmonia, ): Greek goddess of harmony and balance, who brings peace to conflicts. a multi-agent online reinforcement learning (RL) based holistic approach to co-optimize HSS data placement and data migration.
Why RL? We choose RL for data management in HSS because it offers four key benefits compared to heuristic and supervised-learning-based approaches. First, unlike heuristic and supervised learning approaches, RL can learn online and dynamically adapt its decision-making for the current workload and HSS conditions, without prior training using labeled data or human intervention. Second, RL can handle complex tasks such as placement and migration, where each action changes the environment. Third, RL is suitable for even random access patterns by learning the performance impact of actions (i.e., placement or migration) on devices rather than directly predicting future I/O requests. Fourth, RL significantly reduces the system architect’s design effort by offering easy extensibility to various workloads and HSS configurations.
Need for multi-agent RL. We describe the benefits of RL for HSS data management in §3.1. Despite these benefits, a single RL agent cannot effectively (1) co-optimize (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ) different objectives like data placement and data migration, and (2) perform both tasks concurrently. Our motivational study demonstrates that a single-agent RL approach that performs both data placement and data migration leads to 18.8% (27.7%) lower average performance than Sibyl, which performs only data placement, on a performance- (cost-) optimized HSS (see §3.4). This is because learning to optimize data migration interferes with the process of learning an optimal data placement policy as they are two different tasks (See §3.2). A multi-agent RL can mitigate this interference by decoupling the tasks, enabling specialized agents to learn policies that achieve higher HSS performance.
Key Idea. Harmonia employs two coordinating lightweight autonomous RL agents, respectively for data placement and migration, to improve HSS performance. First, we design Harmonia’s data-placement agent to identify the most suitable storage device in the HSS for storing incoming pages. It builds upon the state-of-the-art placement technique, Sibyl (singh2022sibyl, ), but extends it to facilitate coordination with the migration agent. For Harmonia’s data-placement agent, we introduce new migration-aware state features and enhance its reward structure to ensure that placement actions effectively utilize the HSS storage devices. Second, we design Harmonia’s data-migration agent to continuously monitor previously-placed pages in HSS and identify migration candidates and determine their target devices. Harmonia assigns rewards to the migration agent based on the impact of its migrations on future HSS performance. The two agents influence each other’s decisions through actions that affect the HSS environment. This decoupled design prevents policy interference, enables each agent to specialize in its task, and improves overall HSS performance.
We evaluate Harmonia on a real Linux system using seventeen data-intensive workloads and various HSS configurations with up to four devices. Compared to the best-performing prior approach, Sibyl (singh2022sibyl, ), on average, Harmonia provides (1) 49.5% (31.7%) performance improvement on a performance- (cost-) optimized HSS, and (2) 37.0% (42.0%) higher performance on an HSS with three (four) devices. Harmonia incurs low latency (240nsns for inference) and storage overheads (206 KiB for both RL agents together).
This work makes the following key contributions:
  • We show that prior data-management approaches have low HSS performance because they do not optimize both data placement and data migration, and we quantitatively demonstrate their limitations on a real Linux system with various HSS configurations.
  • We propose Harmonia, a multi-agent RL-based data-management technique for HSS that co-optimizes data-placement and data-migration policies using two lightweight coordinating RL agents to improve HSS performance.
  • Via a rigorous real-system evaluation of Harmonia on various HSS configurations, we demonstrate that Harmonia consistently outperforms state-of-the-art data-management techniques across a wide range of data-intensive workloads.
2. Background
2.1. Hybrid Storage Systems
Fig. 1 presents an overview of a hybrid storage system (HSS). An HSS has two key components: (1) multiple storage devices of varying characteristics (e.g., I/O latency, I/O bandwidth, and device capacity) and storage protocols (e.g., NVM Express (NVMe) (nvme, ) or SATA (sata, )), and (2) an HSS management layer that orchestrates data placement and data migration in the HSS.
Figure 1. Overview of a hybrid storage system.
A typical HSS combines fast storage devices with lower capacities (i.e., high-end device in Fig. 1) and slower storage devices with larger capacities (i.e., mid-range and low-end devices in Fig. 1).
HSS Management Layer. The HSS management layer has three key functions. First, it (1) provides a unified flat virtual address space for the applications to store data across all HSS devices, (2) maps a virtual address to a logical block address (LBA) in one of the underlying storage devices, and (3) translates application I/O requests to the devices using NVMe (nvme, ) or SATA (sata, ) protocols. The HSS management layer is similar to the md (mdlinux, ) feature in Linux, which provides software RAID functionality (RAID 0) when multiple storage devices are present. Second, it (1) places data (i.e., data placement) in the best-fit storage device for each incoming I/O request, and (2) rearranges data (i.e., data migration) across storage devices. Prior works (e.g., (singh2022sibyl, ; matsui2017design, ; ren2019archivist, )) demonstrate that data-placement and migration policies have a significant impact on HSS performance. Third, the HSS management layer stores the following metadata in the host DRAM: (1) virtual address to LBA mapping, (2) I/O request characteristics (e.g., request type, request size), and (3) storage device characteristics (e.g., device capacities and their current utilization). The HSS management layer is typically implemented in the kernel space of the host OS.
2.2. Reinforcement Learning
Reinforcement learning (RL) (sutton_2018, ) is a machine learning (ML) approach in which an agent learns the optimal policy for a specific objective by interacting with its environment. Fig. 2 presents the overview of an RL system.
Figure 2. Overview of reinforcement learning.
An RL approach is formulated using four key components.
1. State is the representation of the environment. Assuming $S$ is the set of all possible states, $s_{t}$ $\in$ $S$ is the state at time step $t$.
2. Action. At each time step, the agent observes the current state $s_{t}$ and performs an action $a_{t}$ from the set of all actions $A$. Based on the action, the environment transitions to a new state $s_{t+1}$.
3. Reward. In response to $a_{t}$ that changes environment state from $s_{t}$ to $s_{t+1}$, the agent receives a numerical reward $r_{t+1}$.
4. Policy. An agent uses policy $\pi$ to determine its actions in a given state and aims to find the policy that maximizes cumulative rewards. The optimal policy $\pi^{*}$ is determined by computing the optimal action-value function $Q^{*}$, also called the Q-value. The Q-value of a state-action pair, denoted as $Q(S,A)$, reflects the predicted cumulative reward obtained by taking action A in state S.
The agent makes decisions and learns to maximize the overall cumulative reward or return over time. While the agent may not maximize the benefit of each action, it aims to maximize the long-term consequences of its actions.
2.3. Multi-Agent Reinforcement Learning
Multi-agent RL (e.g., (canese2021multi, ; rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) is an extension of traditional RL where multiple RL agents interact with a shared environment and and with one another to learn optimal policies. Unlike RL with a single agent, multi-agent RL can effectively optimize multiple different tasks. A single RL agent cannot effectively learn optimal policies for different tasks because it relies on task similarity (ammar2014automated, ; carroll2005task, ) to transfer the learning gained from one task to another. Task similarity refers to the resemblance of tasks in their inputs, objectives, and actions. Prior works (yang2020multi, ; yu2020meta, ; teh2017distral, ; sun2020adashare, ; ammar2014automated, ; carroll2005task, ) show that a single RL agent cannot effectively learn optimal policies for multiple tasks with low task similarity because transferring the learning from one task to a dissimilar task can inhibit performance.
3. Motivation
In this section, we describe (1) why RL is a good fit for HSS data management, (2) the need for a multi-agent RL approach, and (3) prior approaches to HSS data management and their effectiveness.
3.1. Why RL for Data Management in HSS?
RL is a good fit for HSS data management for four key reasons.
First, unlike heuristic and supervised-learning approaches, an RL approach can learn online without prior training using labeled data. Online learning helps the RL agent learn dynamic changes in workload access patterns and system conditions, and continuously adapt the data-placement and data-migration policies using system-level feedback without human intervention.
Second, RL can handle complex tasks such as placement and migration where each action changes the environment (e.g., device capacity utilization, bandwidth utilization, garbage collection) and impacts future outcomes (e.g., I/O request latencies). Unlike supervised-learning approaches, RL learns strategies that maximize long-term returns (e.g., (sutton_2018, ; sutton1999policy, ; ke2018modeling, )), i.e., overall HSS performance.
Third, RL is a suitable approach even for random access patterns compared to heuristic and supervised learning because an RL approach learns the expected performance impact of placement or migration on a target storage device instead of predicting future I/O accesses directly. This allows the RL agents to implicitly learn value-based policies, even when the access patterns are random.
To demonstrate the dynamic changes in workload access patterns, we show the virtual addresses accessed and I/O request sizes during a 300-second execution of an example workload, ssd-02 (see §5 for evaluation methodology), in Figs. 3(a) and 3(b) respectively.
Figure 3. Virtual addresses accessed (top) and request sizes (bottom) captured during a 300-second execution of ssd-02 workload. Each virtual address accessed is represented as a green dot in the top figure.
We make two key observations: (1) the application accesses a broad range of virtual addresses ($\times1$$0^{6} - 1$$\times$$10^{9}$) without fixed patterns. (2) the I/O request sizes vary from random accesses (i.e., 4KiB) to large sequential accesses (i.e., 128KiB). We observe such dynamic access pattern variations in all our chosen workloads. HSS system conditions can also change dynamically. For example, the device capacity utilization changes with each write operation, and access latencies vary significantly when multiple applications access a device concurrently. Prior heuristic and supervised learning based approaches (see S3.3) make fixed decisions at infrequent intervals, and cannot adapt efficiently to the dynamic variations in workload access patterns and system conditions (see §3.4 for our motivational results). Prior work Sibyl (singh2022sibyl, ) shows that RL outperforms heuristic-(e.g., (matsui2017design, ; matsui2017tri, )) and supervised-learning-based (e.g.,  (doudali2019kleio, ; shetti2019machine, )) approaches for HSS data placement. However, Sibyl partially uses RL for HSS data-management (i.e., RL for data-placement and a heuristic approach for data migration) that leaves a large gap between its performance and that of an Oracle approach (as we discuss in §3.4).
Fourth, RL offers ease of extensibility to a wide range of HSS configurations and devices with heterogeneous characteristics with minimum designer effort (as partially demonstrated in Sibyl (singh2022sibyl, )).
3.2. Need for Multi-Agent RL for HSS
Data placement and data migration are two different tasks in HSS, and a multi-agent RL approach is a great fit for optimizing these different tasks (see §2.3).
Data placement and data migration in HSS are two different tasks for three key reasons. First, they operate on different data in the HSS. While data placement writes data from the current I/O request to the best-fit storage device, data migration monitors previously placed data to identify candidate pages for migration and their target storage devices. Second, their objectives are different. The objective of data placement is to minimize the I/O latency for each incoming request, but data migration aims to improve long-term HSS performance. Third, these two tasks are performed asynchronously with each other. Data placement is on the critical path of I/O handling and is performed synchronously with application writes or updates. Data migration is a background process and it is performed asynchronously to the application’s I/O requests.
To describe the need for data migration in addition to data placement, we provide two examples: (1) If a page in the fast device is rarely accessed, migrating it to the slow device can free up space for more frequently-accessed pages and improve overall HSS performance. (2) If a page on the slow device is accessed frequently (due to access pattern changes), it should be migrated to the fast device to reduce its access latency. In §3.4, we quantitatively demonstrate that a single RL agent (SAPM) shows limited performance when handling both data placement and migration in HSS. We conclude that multi-agent is a great fit for optimizing both data placement and migration in HSS, which are complex and different tasks.
3.3. Prior Approaches to HSS Data Management
Our focus in this work is to co-optimize placement and migration in HSS. In this section, we briefly introduce prior HSS data-management techniques (see Table 1). However, there is no existing technique that optimizes both placement and migration in HSS. We discuss different approaches to optimize both placement and migration by combining (or extending) prior HSS data-management techniques.
Table 1. Comparison of HSS Data-Management Techniques.
Technique
Data
Placement
Data
Migration
Key Details
CDE (matsui2017design, ) Heuristic Heuristic
Both policies are hand-tuned.
Data eviction during
I/O request handling.
RNN-HSS (doudali2019kleio, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
K-SVM (shetti2019machine, ) Heuristic
Supervised
Learning
Needs frequent retraining.
Data migration at fixed intervals.
Sibyl (singh2022sibyl, ) RL Heuristic
Data eviction during
I/O request handling.
Sibyl (singh2022sibyl, )
+ K-SVM (shetti2019machine, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
Sibyl (singh2022sibyl, )
+ RNN-HSS (doudali2019kleio, )
RL
Supervised
Learning
Lack of coordination
between the two policies.
CDE (matsui2017design, )
+ RL-Migr
Heuristic RL
Hand-tuned data-placement policy.
SAPM RL
Single RL agent for both tasks.
Harmonia RL RL
Two coordinating RL agents.
Data-Placement Techniques. We evaluate two state-of-the-art HSS data-placement techniques: CDE (Cold Data Eviction) (matsui2017design, ) and Sibyl (singh2022sibyl, ). CDE is a heuristic data-placement technique that places hot or random (cold or sequential) data in the fast (slow) device, by determining the hotness and randomness of incoming data based on past access frequency and request size, respectively. Sibyl (singh2022sibyl, ) uses an RL agent to place data in the best-fit HSS storage device based on multiple features of I/O requests (e.g., request type and size) and HSS conditions (e.g., device capacity, I/O latency). When the fast device becomes full, both techniques evict the least-recently-used (LRU) data before placing new I/O request data in the fast device.
Data-Migration Techniques. We evaluate two state-of-the-art HSS data-migration techniques, K-SVM (shetti2019machine, ) and RNN-HSS (doudali2019kleio, ), which have three aspects in common. First, they both leverage supervised learning to identify the target device for migration of each stored page within the HSS based on the page’s access frequency. Second, they periodically determine migration candidates (i.e., pages) and perform data migration at fixed intervals (e.g., migration is performed after every 1000 incoming requests) on the critical path of I/O request handling. Third, they perform data placement for incoming I/O requests using a fixed heuristic policy (e.g., K-SVM places incoming requests in the fast device until its capacity reaches 100% (shetti2019machine, )). The key difference between the two techniques is in the supervised-learning method used; K-SVM uses a k-means-assisted support vector machine (SVM) to classify stored pages as belonging to a storage device based on their past access frequencies; RNN-HSS uses recurrent neural networks (RNN) to predict a page’s future access frequency, and migrate the page to its target device.
Extended Data-Management Techniques. Since there is no existing technique that optimizes both data placement and data migration together, we evaluate four extended techniques to explore the design space for a holistic data-management technique. For the first two techniques, Sibyl+K-SVM and Sibyl+RNN-HSS, we simply combine Sibyl, the current best-performing data-placement technique, with the prior data-migration techniques explained above.
We design the third extended data-management technique, CDE+RL-Migr, [^2] The RL-based data-migration technique in CDE+RL-Migr is a contribution of this work. We design this as there are no prior RL-based HSS data-migration techniques. by combining CDE, a heuristic data-placement technique, with an RL-based data-migration approach. The RL-based data-migration approach uses the same design as the data-migration agent proposed in Harmonia, including the state features and the reward structure (See §4.2).
In the fourth extended technique, SAPM (Single Agent for Placement and Migration), we design a single RL agent to perform both data placement and data migration to analyze the effectiveness of a single RL agent at HSS data management. SAPM performs a placement action for every incoming I/O request, and focuses on data migration only during system idle times to avoid interference with the placement task (which is on the critical path of I/O request handling). We provide the same state features used in Harmonia (See §4.2) to SAPM and tune the hyperparameters to achieve the best performance possible with a single RL agent. We provide SAPM’s RL agent (1) an immediate reward for every placement action (based on the reward of Harmonia’s data-placement agent in Equation 2), and (2) a deferred reward for the migration actions (based on the reward structure of Harmonia’s data-migration agent in Equation 3). We provide background in §2.3 on why a single RL agent cannot effectively optimize two different tasks.
3.4. Effectiveness of Prior Techniques
Methodology. We evaluate eight prior data-management techniques on a real system with two HSS configurations, performance- and cost-optimized, on 17 data-intensive workloads from 3 benchmark suites (see Table 5). See §5 for our evaluation methodology.
Performance Results. Fig. 4 shows the performance of four prior and four extended HSS data-management techniques on (a) performance-optimized and (b) cost-optimized HSS. We compare their average request latencies against two ideal approaches, Fast-Only and Oracle. First, Fast-Only is an ideal policy in which we assume that the fast storage device is large enough to accommodate the entire workload dataset. Second, Oracle (meswani2015heterogeneous, ) makes optimal data-placement and migration decisions based on complete knowledge of future I/O access patterns of the entire workload. We assume that Oracle performs data migration during system idle times, which does not add any latency overheads. Oracle makes data-placement and data-migration decisions while taking the device capacities into account. All values in Fig. 4 are normalized to Fast-Only.
Figure 4. Performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM and Oracle on performance-optimized (top) and cost-optimized (bottom) HSS configurations. Performance is shown as average I/O latency normalized to Fast-Only. Lower is better.
We make six key observations from Fig. 4. First, all prior techniques fall significantly short of Oracle for every HSS and workload. The best-performing technique, Sibyl, achieves only 49.9% (41.4%) of Oracle’s performance in performance- (cost-) optimized HSS, highlighting the need for holistic HSS data management. While Sibyl’s RL-based placement can adapt to workload and system changes, its heuristic eviction on the critical path limits performance.
Second, Sibyl outperforms CDE, K-SVM, and RNN-HSS by 20.7% (14.5%), 24.7% (31.6%), and 19.3% (19.2%) in the performance- (cost-) optimized HSS, respectively, due to its adaptability via RL, unlike heuristic- or supervised-learning-based approaches, which cannot easily adapt to workload changes or HSS conditions.
Third, Sibyl+K-SVM and Sibyl+RNN-HSS underperform compared to Sibyl, as naive combinations of uncoordinated placement and migration policies can lead to (1) conflicing decisions, and (2) slow convergence.
Fourth, CDE+RL-Migr shows lower performance than Sibyl on average (4.6% (15.6%) in performance- (cost-) optimized HSS). It has comparable performance to Sibyl in read-intensive workloads (i.e., RocksDB, MLPerf) due to proactive prefetching of frequently-read data by its RL-based migration agent but suffers in write-intensive workloads (i.e., SYSTOR17) due to its heuristic placement policy, which results in frequent migrations. CDE+RL-Migr’s performance suggests that an RL-based data-migration policy can benefit performance if combined with an adaptive placement policy.
Fifth, SAPM underperforms Sibyl by 18.8% (27.7%) on average in the performance- (cost-) optimized HSS, limited by (1) its single RL agent handling both placement and migration tasks, and (2) inability to identify migration candidates in a timely manner because it focuses on migration only when there are no incoming I/O requests. We provide a detailed analysis on why a single RL agent cannot optimize two different tasks in §3.2.
[…]
3.5. Our Goal
Based on our observations and analyses in §3.1, §3.2, §3.3, and §3.4, we conclude that a multi-agent RL approach’s ability to optimize different tasks concurrently alleviates the limitations of a single-agent RL approach.
Our goal is to design a holistic HSS data-management technique that optimizes both data-placement and data-migration policies using a multi-agent RL approach to improve the overall HSS performance.
4. Harmonia
We design Harmonia as an online multi-agent RL technique for HSS data management with two autonomous RL agents: a data-placement agent and a data-migration agent. We implement Harmonia in the HSS management layer in the host operating system. In this section, we first describe the challenges of designing a multi-agent RL-based technique for HSS data management, and then explain Harmonia’s design and implementation in detail.
4.1. Challenges of Designing a Multi-Agent HSS Technique
Designing a multi-agent RL-based HSS data-management technique poses three key challenges. First, we need to design two autonomous RL agents to optimize the different tasks of data placement and migration (see §3.2). Second, we need to formulate the reward structures of the two RL agents to enable coordination between them (see §4.3) and improve HSS performance. Third, having multiple RL agents can increase computational complexity and storage overhead (see §4.6). While multi-agent RL itself is not a fundamentally new approach to system optimization, prior techniques (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) cannot be adapted directly for HSS data management because they use (1) identical RL agents with a single objective, and (2) joint-action space and shared reward functions for all RL agents.
4.2. Harmonia: RL Formulation
Fig. 5 shows the overview of Harmonia’s multi-agent RL framework. The key goal of the data-placement agent is to place incoming I/O requests in the best-fit storage device, and that of the data-migration agent is to identify migration candidates from previously-placed pages and determine their target storage device. We design the state, action, and reward of the two agents to enable coordination between them and maximize the overall HSS performance.
Figure 5. RL Agents in Harmonia.
State. We design a common set of seven state features because both agents consider the current HSS conditions and data characteristics when making their decisions. Table 2 summarizes the state features for Harmonia’s two RL agents. Request type (req_typereq\_type) indicates a read or a write operation. Request size (req_sizereq\_size) is the number of 4KiB pages in the current I/O request. Access interval (acc_intracc\_intr) is the number of page accesses between two accesses to the same page, while access frequency (acc_freqacc\_freq) is the total number of accesses to the page. Fast-device capacity (fast_capfast\_cap) is the remaining free space in the fast storage device. Current device (curr_devcurr\_dev) indicates the storage device where the page currently resides. Migration interval (migr_intrmigr\_intr) is the number of migrations performed between two migrations of the current page. The state features are represented as a 7-dimensional tuple (observation vector) for both agents:
(1) $O_{t}=(req\_size_{t},req\_type_{t},acc\_intr_{t},acc\_freq_{t},\\ fast\_cap_{t},curr\_dev_{t},migr\_intr_{t}).$
The observation vector size directly impacts Harmonia’s storage and computational overheads (see §4.6), so we (1) carefully select a limited set of highly relevant features, and (2) group state feature values into a small number of bins (Table 2) based on workload characterization. For example, request size can fall into eight size classes (e.g., $\leq$ 4 KiB, 8 KiB, 16 KiB) requiring three bits to represent. In total, we use 32 bits for encoding the state representation. Based on empirical analysis, we use additional bits for some state features (e.g., acc_intr, acc_freq, migr_intr) to represent the values accurately and provide extensibility to new workloads and more storage devices in the future.
Harmonia maintains a metadata table in the host DRAM to track the state features, and updates this table for each I/O request. Harmonia performs a metadata table lookup to observe the current state of a page in the HSS for placement and migration decisions.
Table 2. State features used by Harmonia
Feature Description # of bins Encoding (bits)
req_typereq\_type Request type (read/write) 2 1
req_sizereq\_size Request size (in pages) 8 3
acc_intracc\_intr
Access interval
of the requested page
64 8
acc_freqacc\_freq
Access frequency
of the requested page
64 8
fast_capfast\_cap
Free space
in the fast storage device
8 3
curr_devcurr\_dev
Storage device where the
requested page currently resides
2 1
migr_intrmigr\_intr
Migration interval of a page
64 8
Action. Harmonia’s agents operate independently and their action space depends on the number of devices in the HSS. The data-placement agent selects a device for each incoming write request, while the data-migration agent decides migration targets for existing data. In an HSS with two devices, the agents must choose between fast and slow devices. The small, extensible action sets reduce complexity and support multi-device scalability, with actions encoded with 4 bits to allow extension to more storage devices.
Reward. The reward structure is key to finding an optimal policy in RL. We design the reward structures of Harmonia’s agents based on two key factors. First, since we focus on HSS performance, we use I/O request latencies in our reward structure for both agents. I/O request latencies capture the internal state (e.g., queueing delays, buffer dependencies, bandwidth utilization, effects of garbage collection, read/write latencies, non-blocking operations, error handling latencies) of storage devices in the HSS. Second, we design the reward structures to enable coordination between the two agents.
For every placement action at time step $t$, the data-placement agent receives a reward $R_{placement}$ from the environment at time step $t+1$ that is inversely proportional to the I/O request latency:
(2) $R_{placement}=\frac{1}{L_{t}}$
where $L_{t}$ is I/O request latency at time step $t$. The reward is higher (lower) if the data is placed in the fast (slow) device. We design a simple reward for the data-placement agent because its objective is to minimize the latency of the ongoing I/O request. While placing data on the fast device yields lower latency and a higher reward, Harmonia does not always place data in the faster device for two reasons: (1) Fast device capacity is limited, and overusing the fast device can degrade performance due to factors such as queueing delays and garbage collection. (2) RL agents learn to maximize long-term cumulative reward. Hence, they learn to balance current latency against future access patterns and device utilization. Therefore, Harmonia places data on the slow device to prevent contention, and reserves the fast device for critical data.
The data-migration agent’s objective is to maximize long-term HSS performance, which is different from the placement agent’s objective (see §3.2). Unlike the immediate impact of data placement on HSS performance, the impact of data migrations (i.e., prefetching frequently-accessed data to the fast device and evicting cold data to the slow device) may be delayed due to two key reasons: (1) the application may not access the prefetched data immediately, and (2) the application may not immediately leverage the free space created in the fast device by cold data migration. Hence, we provide a delayed reward to the migration agent to capture the long-term impact of its decisions. The reward for the migration agent is:
[…]
where $t$ is the current time step, $n$ is the number of incoming I/O requests whose data-placement latencies are considered, $L_{i}$ is the latency of I/O request $i$, $P_{migr_{t}}$ is a penalty based on migration and access intervals, and $x$ is the number of pages migrated before providing a reward. This reward is based on the average data-placement latencies (i.e., $\frac{\sum L_{i}}{n}$) for $n$ I/O requests that arrive after $x$ migration candidates are moved to their target devices, and serves as feedback for the migration decisions made previously. To avoid repeated migrations of the same page (ping-pong migrations), we add a small penalty $P_{migr_{t}}$ whose value is inversely proportional to the average migration and access intervals (See Table 2) of $x$ pages. Based on empirical analysis (see §6.1), we set the value of $n$ to 50 and $x$ to 10, the size of the migration queue. We use a half-precision floating-point (16-bit) representation for the reward structure of both agents to reduce the storage overhead of training data.
4.3. Coordination Between Harmonia’s Agents
We achieve coordination between the two agents through (1) their rewards, and (2) the effects of their actions on the HSS environment. The placement agent’s reward (See Equation 2) encourages effective utilization of the fast device, which is possible only if the migration agent proactively migrates cold (hot) data to the slow (fast) device. The data-migration agent receives a delayed reward (See Equation 3) that is based on access latencies of $n$ incoming I/O requests after migrating several candidate pages. In this way, each agent’s decisions influence the policy learned by the other agent.
Mitigating Conflicting Decisions. To prevent conflicting placement and migration decisions, we use two key techniques. First, the migration agent prioritizes the monitoring of pages with high access and migration intervals, and ignores recently placed or migrated pages (e.g., if the page’s access interval is 1, i.e., it was recently placed). Second, we add a small penalty based on access and migration intervals to the migration agent’s reward (See Equation 3) to discourage frequent page movements.
Impact of Coordinated vs. Uncoordinated Agents. To demonstrate the impact of coordination between Harmonia’s placement and migration agents, we compare the performance of Sibyl and Harmonia against Harmonia without coordinated agents in Figs. 6(a) and  6(b) on a performance-optimized and cost-optimized HSS, respectively. We design Harmonia without coordination by devising a reward for the migration agent that improves local performance (i.e., latency cost of page migrations) without considering the placement agent’s actions. While there are other possible reward structures, we choose one that does not facilitate coordination.
Figure 6. Performance of Sibyl, Harmonia without coordination and Harmonia on performance- (cost-) optimized HSS. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Harmonia without coordination by 35.1% (42.1%) on average on a performance- (cost-) optimized HSS. The two agents in Harmonia without coordination are oblivious to each other and take conflicting decisions. Second, Harmonia without coordination performs poorly in read-intensive workloads (e.g., RocksDB) because the migration agent does not prioritize migrating pages to the fast device that will likely be read in the future, which results in lower performance than Sibyl.
4.4. Harmonia: Design
Fig. 7 shows the key components of Harmonia. We describe these key components below.
Figure 7. Overview of Harmonia.
Data-Placement Agent  1. An RL agent that places I/O request data in the best-fit storage device in the HSS 2 based on I/O request characteristics and HSS conditions. This agent consists of two threads. First, the data-placement training thread  3 is a background process in which the agent is trained. Second, the data-placement inference thread  4 is Harmonia’s only foreground thread. In this main decision thread, the data-placement agent decides the target storage device for incoming I/O request data on the critical path.
Data-Migration Agent 5. We implement the data-migration agent using two background threads to avoid interference with the I/O request handling on the critical path. First, the data-migration training thread  6 is used for training the agent. Second, in the data-migration inference thread  7, the agent performs three key tasks: (i) continuously monitors previously-placed pages in the HSS. (ii) determines migration candidates and their target devices. (iii) pushes the metadata of the migration candidates into the migration queue  8. A page is not a migration candidate if the target device determined by the migration agent is the same as its current device.
Migration Queue 8. A queue that stores the logical page address and the target device identifier (e.g., 0 for slow device) of each migration candidate identified by the migration agent. Based on empirical analysis (see §6.1), we set the migration queue size to 10 in our experiments to migrate pages in a timely manner during system idle times and avoid stale migration candidates.
Migration Thread 9. A background process that continuously monitors the migration queue and migrates pages across devices during system idle times. Harmonia samples device bandwidth at frequent intervals using real-time tools (e.g., iostat) to perform migrations. For high-intensity workloads with short system idle times, the migration thread checks if pages related to incoming read requests are already in the migration queue and issues low-priority writes (nvme, ; sata, ) to immediately schedule them for migration.
Harmonia’s five-threaded design prevents the following background processes from interfering with data placement of incoming I/O requests: (1) agent training, (2) migration candidate identification, and (3) data migration across storage devices in the HSS.
4.5. Design of Harmonia’s RL Agents
Fig. 8 shows the detailed design of each RL agent. Both agents in Harmonia have a similar design using two threads (training and inference) for ease of implementation and low storage overhead.
Figure 8. Design of RL Agent in Harmonia.
Inference Thread. Harmonia’s RL agents use the inference thread to make decisions while collecting experiences to use as training data. In this thread, the agent (1) observes the input features (state) a from the I/O request data and HSS, (2) performs inference using its inference network b, (3) selects the action c that maximizes d its long-term returns, and (4) collects the state, action and reward e as experiences f in an experience buffer g.
Training Thread. In this background thread, Harmonia uses collected experiences h from the experience buffer to train the training network i and optimize the placement or migration policy.
Experience Buffer. The two agents collect $\langle State,Action,Reward,NextState\rangle$ tuples (experiences) via their inference threads and store them in separate experience buffers, reflecting their distinct goals. The experience buffers reside in the host DRAM, and we use empirical analysis to set their sizes to store the most recent 1000 experiences. A larger experience buffer can capture more access patterns but increases storage overhead. Harmonia uses experience replay (dqn, ; singh2022sibyl, ) to sample a batch of experiences from the experience buffers to train its agents.
Training and Inference networks. We separate the RL agent’s training and inference networks to remove training latency from I/O request handling. Weights are copied from training network to inference network every 1000 requests to keep inference up to date. Both agents use identical simple feed-forward networks (singh2022sibyl, ) with one hidden layer of 10 neurons (de1993backpropagation, ) and swish activation (ramachandran2017searching, ). Actions are selected based on the highest predicted Q-value.
RL Algorithm. We use C51, a categorical Deep Q-Network (C51, ) for both data-placement and data-migration agents to update the Q-values $Q(S,A)$ (See §2). C51 aims to learn the distribution of Q-values for each state-action pair, which helps Harmonia capture more information from the environment (harrold2022data, ; singh2022sibyl, ).
Hyper-Parameter Tuning. We improve Harmonia’s accuracy by tuning the hyper-parameters (paine2020hyperparameter, ) of each agent using cross-validation (arlot2010survey, ) of different values. Harmonia’s agents need only one-time offline hyper-parameter tuning. Table 3 shows the hyper-parameter values chosen based on empirical analysis.
Table 3. Hyper-parameters
Hyper Parameter
Design Space
Placement
Agent
Migration
Agent
Discount Factor
($\gamma$)
0-1 0.9 0.1
Learning Rate
($\alpha$)
$1e^{-5}-1e^{0}$ $1e^{-3}$ $1e^{-2}$
Exploration Rate
($\epsilon$)
0-1 0.001 0.001
Batch Size
64-256 128 256
Experience Buffer Size
10-10000 1000 1000
The discount factor ($\gamma$) determines the balance between immediate and future rewards, while the learning rate ($\alpha$) is the rate at which neural network weights are updated. The exploration rate ($\epsilon$) balances exploration and exploitation for Harmonia’s policies. Batch size determines the number of samples processed in each training iteration. Experience buffer size is the number of most recent experiences stored for training the networks.
Exploration vs. Exploitation. Harmonia’s RL agents start with no prior knowledge of the workload or HSS, and make random initial decisions (explore), and use their experiences (exploit) to gradually make optimal decisions. To balance exploration and exploitation, we use the $\epsilon$-greedy policy (tokic2011value, ): the predicted best action is selected with (1-$\epsilon$) probability, and a random action with $\epsilon$ probability.
Convergence of RL Agents in Harmonia. Figs 9(a) and  9(b) show the convergence of Harmonia’s data-placement and data-migration policies in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 5) on a performance-optimized HSS. Both agents start with no prior training (i.e., cold start).
Figure 9. Convergence of Harmonia’s (a) Data-Placement and (b) Data-Migration RL agents shown in terms of reduction in training loss. We show only the first 10000 I/O requests.
We make two key observations: (1) The placement policy converges in less than 6000 I/O requests. (2) The migration policy takes longer to converge (around 9000 I/O requests) due to indirect and delayed reward to the migration agent. The training loss remains low after convergence during the workload lifetime. After the initial convergence, Harmonia’s agents are well-equipped to adapt to new access patterns very quickly. They no longer need 6000 requests to start making good decisions when access patterns change.
Adaptability to Workload Changes. To demonstrate Harmonia’s adaptability to changes in workload access patterns, we design a synthetic benchmark using two workloads, LUN0 (write-intensive) and ResNet50 (read-intensive) (see Table 5), alternately every 5000 requests. Figs 10(a) and  10(b) show the access pattern variations in our benchmark in terms of virtual addresses accessed and request sizes. Fig 10(c) shows Harmonia’s convergence in terms of training loss for the first 20000 requests on a performance-optimized HSS. While the access patterns change significantly at every 5000 requests, each workload phase has considerable access pattern variations.
Figure 10. Workload access patterns represented in terms of (a) virtual addresses accessed, and (b) request sizes. Convergence of Harmonia shown in terms of training loss (c) for the first 20000 I/O requests. Lower is better.
Both agents start with no prior training (i.e., cold start). We make two key observations: (1) The rapid decline in training loss at the start indicates that both agents learn the optimal policy from a cold start in less than 5000 requests (Phase 1: LUN0). (2) When the access patterns change significantly after 5000 requests (Phase 2: ResNet50), the training loss remains low (nearly 0). (3) We observe a minor spike (shown in the inset plot) in training loss at every phase change. However, the training loss is still low, and Harmonia’s policies quickly converge again (in a few hundred I/O requests).
We conclude that Harmonia adapts quickly to abrupt changes in workload after the initial convergence.
4.6. Overhead Analysis
In this section, we analyse Harmonia’s overhead in terms of (1) storage overhead, and (2) inference and training latencies.
4.6.1. Storage Overhead.
[…]
4.6.2. Latency Overhead.
Harmonia has low inference and training latencies because of its (1) lightweight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). We perform training and inference on the host CPU as the neural networks are lightweight, and their weights fit in the CPU on-chip caches in our evaluated system.
[…]
We conclude that Harmonia’s overheads are justified for the performance improvement it achieves for two key reasons. First, Harmonia’s RL agents have a significantly lower memory footprint compared to the that of best-performing prior approach, Sibyl. While Sibyl’s RL agent consists of 58 neurons, Harmonia’s RL agent is light-weight with only 19 neurons. Harmonia’s fewer neurons results in lower inference and training latency compared to Sibyl. Second, Harmonia provides a significant performance improvement (49.5% (31.7%) on performance- (cost-)optimized HSS) over Sibyl.
4.7. Architectural Differences over Prior Works
We describe Harmonia’s novelty and four key architectural differences compared to prior RL techniques (e.g., Sibyl) for HSS.
Separation of Objectives via Multi-Agent RL. Unlike Sibyl’s single-agent RL design, Harmonia decouples placement and migration into two specialized agents, which allows them to learn distinct policies: the placement agent focuses on immediate performance, while the migration agent optimizes long-term benefits. This design improves adaptability in dynamic workload and HSS conditions.
Proactive Data Migration. Harmonia learns to proactively migrate data (even in read-intensive workloads), rather than reacting only when the fast device capacity reaches 100% (as Sibyl does via eviction) or when there are updates from the application. The migration agent optimizes long-term HSS performance resulting in better, more robust handling of dynamic workloads.
Non-trivial RL Agent Formulation. Although Harmonia’s RL agent formulation shares some aspects (e.g., input features, latency-based reward) with Sibyl, it differs in three ways: (1) Harmonia’s neural networks are significantly more light-weight (90% fewer neurons), (2) The input features include migration characteristics (e.g., migration interval) to avoid ping-pong migrations, and (3) Harmonia’s reward structures are designed to optimize performance while facilitating coordination between the agents (See §4.3).
Extensibility. Harmonia’s multi-agent architecture is more amenable to further extensions such as integrating prefetching or load balancing agents.
5. Evaluation Methodology
We evaluate Harmonia on a real system with an HSS with two (dual-HSS), three (tri-HSS), and four (quad-HSS) devices. The dual-HSS configurations include: (1) performance-optimized HSS, which combines a high-end SSD (inteloptane, ) with a mid-range SSD (intels4510, ), and (2) cost-optimized HSS, which pairs the high-end SSD (inteloptane, ) with a low-end HDD (seagate, ). We implement a custom block device driver in Linux for Harmonia to interface with HSS devices. We implement Harmonia using the TF-Agents (TFAgents, ) library. Table 4 describes our real system and the HSS configurations.
Table 4. Host System and HSS Configurations
Host System
AMD Ryzen 7 2700G(amdryzen, ),
[…]
4 MiB L2, 8 MiB L3,
16 GiB RDIMM DDR4 2666 MHz
Storage Devices Characteristics
H: Intel Optane SSD P4800X (inteloptane, )
375 GB, PCIe 3.0 NVMe, SLC,
R/W: 2.4/2 GB/s,
random R/W: 550000/500000 IOPS
M: Intel SSD D3-S4510 (intels4510, )
1.92 TB, SATA TLC (3D),
R/W: 560/510 MB/s,
random R/W: 895000/21000 IOPS
L: Seagate HDD ST1000DM010 (seagate, )
1 TB, SATA 6Gb/s 7200 RPM
Max. Sustained Transfer Rate: 210 MB/s
LSSD: ADATA SU630 SSD  (adatasu630, )
960 GB, SATA, TLC,
R/W: 520/450 MB/s
PMEM (Emulated): Intel Optane
Persistent Memory 200 Series  (intelpmem200, )
128 GB, Memory Mode
R/W: 7.45/2.25 GB/s (256B)
HSS Configurations Devices
Performance-Optimized high-end (H) & middle-end (M)
Cost-Optimized high-end (H) & low-end (L)
HSS with PMEM Emulated PMEM (PMEM) & high-end SSD (H)
Tri-HSS
high-end (H) & middle-end (M) & low-end (L)
Quad-HSS
high-end (H) & middle-end (M) &
low-end SSD (LSSD) & low-end HDD (L)
Baselines. We compare Harmonia against four state-of-the-art heuristic- and machine-learning-based data-placement and migration techniques, (1) CDE (matsui2017design, ), (2) RNN-HSS (adapted from  (doudali2019kleio, )), (3) K-SVM (shetti2019machine, ), and (4) Sibyl (singh2022sibyl, ). We evaluate four extended approaches, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM to explore the design space of holistic data-management techniques. We also compare against two ideal approaches: (1) Fast-Only, and (2) Oracle (meswani2015heterogeneous, ) (See § 3). Oracle serves as a reference to evaluate the accuracy of Harmonia’s decisions as it makes optimal decisions based on knowledge of future access patterns.
Workloads. We select seventeen data-intensive storage workloads from SYSTOR ’17 (lee2017understanding, ), RocksDB traces (yadgar2021ssd, ), Yahoo! Cloud Serving Benchmark (YCSB) suite (cooper2010benchmarking, ), and MLPerf Storage [^3] We generated the traces by running MLPerf applications on our evaluated system. (mlperfstoragebenchmarks, ) from real enterprise and datacenter environments. The average workload size is approximately 50000x the fast device capacity. We choose these workloads to represent diverse I/O access patterns with different read and write ratios, I/O request sizes, and inter-request times. Table 5 reports the characteristics of the chosen workloads. In our evaluation, each workload runs in a separate thread.
Table 5. Characteristics of the evaluated I/O traces
Benchmark
Suite
Traces Read %
Avg. Request
Size (KB)
Avg. Inter-
[…]
SYSTOR17  (lee2017understanding, ) LUN0 0.2 31.7 1163.9
LUN1 0.3 34.2 1864.1
LUN2 7.6 31.1 1418.9
LUN3 3.5 42.7 734.5
LUN4 0.5 26.3 823.1
RocksDB  (yadgar2021ssd, ) ssd-00 79.9 108.9 66.4
ssd-01 73.5 75.1 40.7
ssd-02 79.9 7.5 3.3
ssd-03 79.9 9.5 3.5
ssd-04 79.9 7.8 3.6
YCSB  (cooper2010benchmarking, ) YCSB-B 51.3 45.9 9.3
YCSB-C 47.6 54.6 6.5
YCSB-D 55.9 36.1 8.5
YCSB-E 52.1 46.6 9.6
YCSB-F 49.5 53.1 6.6
MLPerf Storage  (mlperfstoragebenchmarks, ) ResNet50 80.0 172.6 500.1
CosmoFlow 83.4 180.1 1023.8
To evaluate Harmonia under real-world scenarios, we generate six multi-programmed workloads by running multiple workloads concurrently (see Table 6). We choose these multi-programmed workloads based on three key factors: (1) combination of read-intensive, write-intensive and mixed workloads, (2) number of concurrent workloads required for high I/O intensity (i.e., shorter inter-request times), and (3) combination of different request sizes. Each constituent workload is executed using a single thread. For example, mix6 runs on eight concurrent threads for the eight constituent workloads. Multi-programmed workloads typically exhibit higher I/O request intensity and varied access patterns.
Table 6. Characteristics of multi-programmed workloads
Mix
Constituent
Workloads (cooper2010benchmarking, ; lee2017understanding, ; yadgar2021ssd, )
Description
mix1 ssd-02 and LUN4
ssd-02 is read-intensive
and LUN4 is write-intensive
mix2
LUN1 and ssd-04
LUN0 is write-intensive
and ssd-04 is read-intensive
mix3
YCSB-C and YCSB-F
Both have near-equal read-write ratio
mix4
ssd-00, ssd-04,
YCSB-A and LUN0
Two read-intensive
and two write-intensive workloads
mix5
ssd-00, LUN0,
YCSB-C and YCSB-F
Read-intensive, write-intensive and
two workloads with a
near-equal read-write ratio
mix6
YCSB-B, YCSB-D, LUN0, LUN1,
LUN4, ssd-00, ssd-02, ssd03
Two with near-equal read-write ratio,
three write-intensive
and three read-intensive
6. Evaluation
6.1. Performance Analysis
Average I/O Request Latency and End-to-End Throughput. Figs. 11(a) and  11(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance- and cost-optimized HSS configurations. Figs. 12(a) and 12(b) show their end-to-end throughput on performance- and cost-optimized HSS, respectively. Performance and throughput are shown in terms of average request latency and IOPS, respectively, normalized to that of Fast-Only.
Figure 11. Performance of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS configurations, shown as average request latency normalized to Fast-Only. Lower is better.
Figure 12. End-to-end throughput of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS. Throughput values (in IOPS) are normalized to Fast-Only policy. Higher is better.
We make four key observations. First, Harmonia consistently outperforms all baselines in both HSS configurations. In performance- (cost-) optimized HSS, Harmonia improves performance by 49.5% (31.7%) and end-to-end throughput by 49.4% (156.2%) over the best-performing prior approach, Sibyl, on average.
Second, Harmonia bridges the performance gap between Sibyl and Oracle by 64.2% (64.3%) and achieves 71.7% (62.2%) of Oracle’s throughput on average in performance- (cost-) optimized HSS.
Third, Harmonia shows higher performance and throughput benefits over baselines in read-intensive workloads (e.g., RocksDB, MLPerf) compared to write-intensive workloads. For example, Harmonia outperforms Sibyl by 57.5% (43.2%) in read-intensive workloads and 41.0% (39.9%) in write-intensive workloads. In read-intensive workloads, unlike Sibyl, Harmonia proactively prefetches frequently-read pages to the fast device to improve read performance even when there are no writes from the application, resulting in lower latencies and higher throughput. In write-intensive workloads, while both Harmonia and Sibyl use RL-based placement, Harmonia’s migration agent, in coordination with the placement agent, further improves performance by frequently migrating cold data from the fast device to accommodate new incoming requests.
Fourth, Harmonia shows higher performance and throughput gains over baselines (other than Sibyl) in cost-optimized than in performance-optimized HSS. This is because (1) data migration in cost-optimized HSS is more expensive than in performance-optimized HSS due to the large latency gap between the two devices, (2) prior approaches perform migration during I/O request handling, which adds significant latency overheads, and (3) Harmonia frees up space in the fast device with proactive migrations and performing migrations during system idle times. Sibyl’s RL agent places only performance-critical data in the fast device, reducing the number of data evictions on the critical path.
Tail Latency. Figs. 13(a) and 13(b) show the 99th and 99.99th percentile I/O latencies (tail latency) of Harmonia and baselines on performance-optimized HSS.
Figure 13. Comparison of tail latencies of Harmonia and baselines in the 99th percentile (top) and 99.99th percentile (bottom) of I/O requests from three representative workloads, LUN0 (write-intensive), ssd-00 (read-intensive) and YCSB-B (mixed), on performance-optimized HSS.
For this analysis, we choose three representative workloads, LUN0 (write-intensive), ssd-00 (read-intensive) and YCSB-B (mixed) (see §5 for workload characteristics).
We make two key observations. First, Harmonia significantly reduces tail latency compared to all baselines across the three workloads. Harmonia reduces the 99th (99.99th) percentile I/O latency by 32.3% (25.3%) on average over Sibyl. Harmonia has lower tail latency because, unlike prior techniques, Harmonia performs data migration during system idle times to avoid interference with ongoing I/O requests. Second, K-SVM, RNN-HSS, Sibyl+K-SVM and Sibyl+RNN-HSS show very high tail latencies because they perform a large volume of migration at fixed intervals, which adds a significant latency overhead to the ongoing I/O requests.
We conclude that Harmonia improves average I/O latency, end-to-end throughput, and tail latency across all workloads due to its holistic and coordinated data-management approach.
Performance Sensitivity to Migration Queue Size. Fig.  14(a) shows the impact of migration queue size on HSS performance.
We make three key observations. First, optimal migration queue size is 10 for our evaluated workloads. This may vary slightly across workloads and HSS configurations. Second, low queue depths (i.e., ¡ 5) may result in many migration candidates being dropped, which leads to performance degradation. If the migration queue size is 0, Harmonia shows poor performance because it does not perform any data migration. Third, larger-sized queues can hold more migration candidates, but the oldest candidates may become stale (i.e., past migration decisions may not correlate with current access patterns and HSS conditions), which adversely impacts HSS performance.
Performance Sensitivity to Migration Agent’s Reward. Fig.  14(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3).
Figure 14. Performance sensitivity to (a) migration queue sizes, and (b) number of future incoming requests used in migration agent’s reward structure. Performance is shown as average I/O latency normalized to Fast-Only. Lower is better.
We make three key observations. First, the optimal number of requests is 50 in our evaluated configuration, but may vary slightly across workloads and HSS configurations, and depends on factors including (1) inter-request times of incoming I/O requests after migration, (2) workload access patterns (e.g., access frequency and reuse distance of a page), and (3) system conditions (e.g., device capacity utilization). Second, considering fewer requests (e.g., ¡ 10) may not provide the right reward to the agent because the effect of migrations is not seen immediately in the placement latencies of these few requests. Third, if we consider more I/O requests (e.g., 250), we see lower performance because of two reasons: (i) the impact of data migration is diluted if we consider a large number of incoming requests, and (ii) I/O requests that arrive too far into the future after data migration may not see a direct impact.
Performance on HSS with Persistent Memory (PMEM) Devices. To demonstrate the Harmonia’s extensibility to emerging persistent memory technologies, we evaluate Harmonia and baselines on an HSS with a persistent memory device (intelpmem200, ) and a high-end SSD (inteloptane, ) (see Table 4). We emulate the PMEM device using RAMDisk because these emerging devices are not easily available. Fig. 15 shows the performance of Harmonia and baselines on HSS with PMEM. Performance is shown as average request latency normalized to Fast-Only (i.e., PMEM device can store all data).
Figure 15. Performance of Harmonia and baselines on an HSS with PMEM (intelpmem200, ) and high-end SSD (inteloptane, ) shown as average request latency normalized to Fast-Only. Lower is better.
We make three key observations. First, Harmonia outperforms the best-performing baseline, CDE, by 15.6% and achieves 80% of Oracle’s performance. Second, CDE outperforms other baselines in this HSS configuration as it places most data in the low-latency PMEM device. However, CDE’s frequent data evictions on the critical path lower its performance compared to Harmonia. CDE performs better in read-intensive workloads because they consist of fewer writes and require fewer evictions. Third, in read-intensive workloads (i.e., RocksDB, MLPerf), Harmonia shows a higher performance improvement as it proactively prefetches data to the PMEM device, which leads to more reads from the PMEM device. We discuss Harmonia’s applicability to emerging memory technologies in §7.
6.2. Multi-Programmed Workloads
Figs. 16(a) and 16(b) show the performance of Harmonia and baselines on multi-programmed workloads (see Table 6) on performance- and cost-optimized HSS, respectively.
Figure 16. Performance of Harmonia and baselines on multi-programmed workloads on performance-optimized (top) and cost-optimized (bottom) HSS. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms prior approaches in all multi-programmed workloads across both HSS configurations. In performance- (cost-) optimized HSS, Harmonia (i) outperforms Sibyl by 32.8% (25.3%), and (ii) achieves 75% (46%) of Oracle’s performance. Second, Harmonia’s performance improves as the number of concurrent workloads increases. On eight concurrent workloads (e.g., mix6), Harmonia shows higher performance gains, 39.4% (41.8%), over Sibyl compared to other multi-programmed workloads. Harmonia’s RL-based policies learn access pattern variations faster than prior approaches.
6.3. Extensibility to Multiple Devices in HSS
We evaluate Harmonia’s extensibility to different HSS configurations by comparing its performance with the best-performing prior approach, Sibyl, on an HSS with three and four devices (see Table 4). We do not evaluate other baselines as they are designed for fixed HSS configurations and lack easy extensibility. Figs. 17(a) and 17(b) show the performance of Sibyl and Harmonia on a Tri- and Quad-HSS. We discuss Harmonia’s extensibility in §7.
Figure 17. Performance of Sibyl and Harmonia on Tri-HSS (top) and Quad-HSS (bottom). Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make two key observations. First, Harmonia outperforms Sibyl by 37.0% (42.0%) on average in Tri-HSS (Quad-HSS) as Harmonia places performance-critical data in the high-end and mid-range devices and uses its migration policy to move cold data to slow devices during idle times. In contrast, Sibyl frequently evicts data to the low-end devices on the critical path, adding significant latency overhead. Second, in read-intensive workloads (e.g., RocksDB, MLPerf), Harmonia shows higher performance gains (up to 58.3%) over Sibyl because Sibyl lacks a migration policy and relies on application updates to migrate data across devices.
6.4. Impact on Device Lifetimes
We evaluate the impact of Harmonia and baselines on storage device lifetimes by measuring their write amplification (WA). WA is the ratio of total data written (including migrations) in the HSS to data originating from the workload. A higher WA adversely impacts device lifetimes. Figs. 18(a) and 18(b) show WA of Harmonia and baselines on performance- and cost-optimized HSS, respectively.
Figure 18. Write amplification of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS. Lower is better.
We make two key observations. First, only Sibyl and RNN-HSS have a lower WA than Harmonia. Harmonia’s average WA of 1.54 (1.55) is higher than Sibyl’s 1.49 (1.45) and RNN-HSS’s 1.08 (1.06) in performance- (cost-) optimized HSS as it proactively prefetches data to the fast device in read-intensive workloads to improve read performance. Second, in both HSS configurations, Harmonia has lower WA (1.24) in write-intensive workloads (i.e., SYSTOR17) compared to its WA (1.8) in read-intensive workloads (i.e., RocksDB, MLPerf). In write-intensive workloads, Harmonia learns to move data during update operations and performs fewer migrations.
Write Traffic Distribution in Harmonia. Fig.  19 shows the write distribution of Harmonia’s migration policy across devices in performance- and cost-optimized HSS.
Figure 19. Write traffic due to Harmonia’s migrations on performance-optimized (left) and cost-optimized (right) HSS.
We make two observations. First, Harmonia migrates more data to slow device as it effectively utilizes fast device for frequently-accessed data. Second, Harmonia has more writes to fast device in read-intensive workloads (i.e., RocksDB, MLPerf) as it proactively prefetches frequently-read data to the fast device even when there are no application updates.
We conclude that Harmonia’s adaptive policies provide higher performance at slightly higher average WA compared to Sibyl.
6.5. Performance on Unseen Workloads
We evaluate Harmonia’s performance on five unseen data-intensive workloads that are not used for hyperparameter tuning. We choose these workloads (wong2024baleen, ; berg2020cachelib, ; pan2021facebook, ) that use a flash cache in front of a bulk storage system from Meta’s production environment. We compare Harmonia to the best-performing prior approach, Sibyl, in Figs.  20(a) and 20(b) on a performance- (cost-) optimized HSS.
Figure 20. Performance of Sibyl and Harmonia on performance-optimized (top) and cost-optimized (bottom) HSS. Performance is shown in terms of average request latency normalized to Fast-Only. Lower is better.
We make two key observations. Harmonia outperforms Sibyl by 31.2% (31.5%) on average on performance- (cost-) optimized HSS. Second, while Harmonia’s hyperparameters may not be optimized for these workloads, its online-learning-based data-placement and data-migration policies adapt to changing workload access patterns better than Sibyl. We conclude that Harmonia’s generality and multi-agent RL framework provide consistent performance benefits across all unseen workloads.
7. Discussion
Alternate Objectives and Reward Structures. Harmonia focuses on improving HSS performance using RL agents with I/O latency-based rewards. However, it can be adapted for other objectives such as device lifetime, fairness, quality of service (QoS), or energy efficiency. For example, optimizing device lifetime may involve using write counts, endurance, write amplification, and device reliability (e.g., NAND type, ECC) in the input features and reward structures. Fairness and QoS can be addressed by incorporating metrics such as bandwidth, queue depth, and application requirements. We leave the exploration of these objectives and rewards to future work.
Extensibility of Harmonia. Harmonia is easily extensible to a large number of storage devices with minimal designer effort. Adding a new device to the HSS requires adding (1) its capacity utilization to the state features, which results in an additional input node per RL agent neural network, and (2) one output node per RL agent neural network. First, the capacity utilization of the new device consumes 3 bits of storage overhead (see Table 2). Second, the additional neural network nodes (1) add 1.3 KiB of storage overhead, and (2) increase inference and training latency by 30–40 $ns$. While Harmonia outperforms prior approaches, scaling to more devices leads to two challenges: (1) larger state and action spaces can delay policy convergence, and (2) an increase in metadata overhead.
Applicability to Other Systems. Harmonia can be extended to other heterogeneous systems such as tiered storage, hybrid memory, disaggregated storage and networked storage. In tiered storage, placement can be heuristic (always place in the top tier), while migration becomes key, potentially requiring multiple agents for each pair of adjacent tiers. Hybrid memory systems may require Harmonia to be redesigned for low-latency decision-making. Disaggregated systems that use CXL (maruf2023tpp, ; gouk2023memory, ; li2023pond, ) can benefit from deploying Harmonia on the CXL fabric controller. We hope Harmonia can aid in designing adaptive policies across diverse architectures.
8. Related Work
To our knowledge, Harmonia is the first work to use multi-agent RL for combined optimization of data-placement and data-migration policies in HSS. In this section, we review other related work in HSS data management, and RL in storage/memory systems.
HSS Data Management. A large body of prior work proposes heuristic- and machine-learning-based data-placement (e.g., (matsui2017design, ; sun2013high, ; heuristics_hyrbid_hystor_sc_2011, ; lv2013probabilistic, ; guerra2011cost, ; elnably2012efficient, ; heuristics_usenix_2014, ; bu2012optimization, ; krish2016efficient, ; tai2015sla, ; zhang2010automated, ; wu2012data, ; iliadis2015exaplan, ; lv2013hotness, ; matsui2017tri, ; feng2014hdstore, ; yang2017autotiering, ; singh2022sibyl, ; doudali2019kleio, ; ren2019archivist, ; cheng2019optimizing, ; shetti2019machine, ; li2017utility_HMM, ; agarwal2015page_HMM, ; agarwal2017thermostat_HMM, ; ham2013disintegrated_HMM, ; lin2011hot, ; salkhordeh2015operating, ; pavlovic2013data, ; meza2013case, ; chiachenchou2015batman, ; kim2011hybridstore, ; wang2019panthera, ; ramos2011page, ; liu2019hierarchical, ; luo2020optimal, ; doudali2021cori, ; sen2019machine, )) and data-migration (e.g., (lin2011hot, ; lin2017efficient, ; lin2018buffer, ; zhang2010adaptive, ; cheng2015amc, ; shetti2019machine, ; vengerov2008reinforcement, ; vasilakis2020hybrid2, )) techniques. The Linux kernel offers utilities with heuristic solutions to manage multiple storage drives. For example, mdadm provides software RAID in the kernel, combining physical disks into a logical volume, but the burden of using mdadm for data placement and migration lies with applications. Similarly, LVM enables virtual partitions across multiple drives, allowing system administrators to resize them and move data across them. In contrast, Harmonia offers adaptive, application-transparent data placement and migration. For data migration, Vengerov (vengerov2008reinforcement, ) applies fuzzy rules in hierarchical storage and Hot Random Off-loading (lin2011hot, ) uses a 0-1 knapsack model. However, neither (1) conducts real-system evaluation or (2) jointly optimizes placement and migration with multiple RL agents.
RL in Storage Systems. Several works (e.g., (liu2019learning, ; yoo2020reinforcement, ; wang2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2023rlalloc, ; li2020mitigating, ; akgun2021machine, ; kang2017reinforcement, ; kang2019q, ; wei2023reinforcement, ; li2024page, ; liaw2024reinforcement, ; wu2018maximizing, ; ha2024rl, )) apply RL to storage systems, focusing on cloud utilization (liu2019learning, ; wang2020reinforcement, ), SSD throughput and QoS (li2023rlalloc, ), and maintenance tasks such as garbage collection and error mitigation (yoo2020reinforcement, ; rl_GC_TECS, ; kang2018dynamic, ; li2020mitigating, ). Harmonia is orthogonal to these works and can be integrated with them in HSS.
RL for Architectural Optimizations. Prior works use RL for data prefetching (e.g., (pythia, ; peled2015semantic, )), memory scheduling (e.g., (ipek2008self, ; multi_scheduler_HPCA_2012, )), cache replacement strategies (liu2020imitation, ) and network-on-chip arbitration (rl_NOC_AIDArc_2018, ; lin2020deepNOC, ). These works provide evidence that RL holds significant potential for intelligent high-performance computer systems.
Multi-Agent RL Systems. Several works (e.g., (rl_opti_TACO_2017, ; qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mao2023multi, ; mlsys2023qiu, ; lu2022rlrp, ; jain2017cooperative, ; jain2017coordinated, ; shen2022multiagent, )) use multi-agent RL (MARL) for system optimization. For example, Qiu et al. (qiu2022reinforcement, ; qiu2022simppo, ; qiu2020firm, ; mlsys2023qiu, ) target FaaS latency, Jain et al. (rl_opti_TACO_2017, ; jain2017cooperative, ) optimize CPU caches, and Shen et al. (shen2022multiagent, ) address cache cleaning in SMR drives. These works use identical RL agents, joint actions, shared rewards, and table-based RL unlike Harmonia’s heterogeneous, autonomous, neural network-based agents. RLRP (lu2022rlrp, ) proposes a deep RL-based replica placement and migration technique for distributed storage systems. RLRP has five key differences from Harmonia: (1) RLRP does not handle typical I/O requests but focuses on replica placement, (2) RLRP’s placement and migration occur at sparse intervals, making it unsuitable for HSS, (3) RLRP retrains its policies only when storage nodes are added or removed, while Harmonia continuously adapts its policies, (4) RLRP considers only a few state features (e.g., capacity utilization), and (5) RLRP’s design has higher storage overhead than Harmonia. We conclude that these reasons make RLRP unsuitable for data placement and migration in HSS.
9. Conclusion
We propose Harmonia, the first multi-agent reinforcement learning based holistic data-management technique for HSS that optimizes data placement and data migration together. We show that Harmonia’s design enables coordination between the data-placement and data-migration agents to improve HSS performance. Our real-system evaluation demonstrates that Harmonia significantly outperforms prior HSS data-management techniques on four HSS configurations across a wide range of workloads. Harmonia is easily extensible to different HSS configurations. The performance benefits of Harmonia come with low latency and storage overheads.